### v1

# Intermediate Relation Size Bounds for Select-Project-Join Query Plans: Asymptotically Tight Characterizations

Abstract.
We study the problem of statically optimizing select-project-join (SPJ) plans where unary key constraints are allowed. A natural measure of a plan, which we call the output degree and which has been studied previously, is the minimum degree of a polynomial bounding the plan’s output relation, as a function of the input database’s maximum relation size. This measure is, by definition, invariant under passing from a plan to another plan that is semantically equivalent to the first. In this article, we consider a plan measure which we call the intermediate degree; this measure is defined to be the minimum degree bounding the size of all intermediate relations computed during a plan’s execution — again, as a function of the input database’s maximum relation size. We present an algorithm that, given an SPJ plan $q$ and a set $\Sigma$ of unary keys, computes an SPJ plan $q^{\prime}$ that is semantically equivalent to $q$ (over databases satisfying $\Sigma$) and that has the minimum intermediate degree over all such semantically equivalent plans. For the types of plans considered, we thus obtain a complete and effective understanding of intermediate degree.
Select-Project-Join Plans, Query Optimization, Tree Decompositions, Key Dependencies
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
Modern database systems capable of evaluating queries of the standard query language SQL typically have, as a key component, a query optimizer that synthesizes execution plans consisting of basic operations from a relational algebra; techniques for estimating and comparing the costs of evaluating such plans are considered central in the study of databases (Abiteboul et al., 1995, Chapter 6). In this article, we study a basic measure of execution plans: the intermediate size, by which we mean the maximum size over all intermediate relations computed during a query’s execution, relative to a database. We view this as a natural and fundamental measure: the intermediate size gives an obvious lower bound on the amount of time needed to execute a plan, for algorithms that explicitly materialize all intermediate relations; and, it characterizes the amount of space needed to store the results of all such intermediate relations.
We study this size measure in an asymptotic setting, as a function of the maximum relation size $M_{\mathbf{D}}$ of a database $\mathbf{D}$. We say that a plan has intermediate degree $d$ when $d\geq 0$ and there exists a degree $d$ polynomial upper bounding the plan’s intermediate size, as a function of $M_{\mathbf{D}}$, but no lower degree polynomial provides such a bound. We study this measure for SPJ query plans, which allow select, project, and join as operations on relations. These plans are intensely studied throughout database theory. We permit unary key constraints to be expressed on database relations. In essence, a unary key constraint states that there is a single coordinate of a relation where, for each tuple in the relation, knowing the value at the coordinate determines the entire tuple.
Our main theorem presents an algorithm that, given an SPJ plan $p$ with a set $\Sigma$ of unary key constraints, computes an SPJ plan $p^{\prime}$ that is semantically equivalent to $p$ (with respect to databases satisfying the constraints $\Sigma$), and has the lowest possible intermediate degree over all SPJ plans that are so semantically equivalent to $p$. This theorem thus yields, in the asymptotic setting considered, a full understanding of how to statically optimize the intermediate degree so that it is the minimum possible. We remark that (to our knowledge) this type of understanding was not previously known even for SPJ plans on databases without any constraints. We view the definition of the asymptotic notion of intermediate degree, and its supporting theory, as conceptual contributions of this work.
Our result showing how to minimize the intermediate degree amounts to demonstrating how to compute, for any given plan, an asymptotically optimal algorithm in a restricted model of computation, namely, the class of SPJ plans. Our focus on polynomial degree is aligned with the notion of fixed-parameter tractability, whereby a problem is tractable essentially when there is a degree $d$ such that each so-called slice has an algorithm with running time bounded above by a degree $d$ polynomial; a slice is a set of all instances sharing the same parameter, which is a value associated with each instance.
To prove our result, we make a number of technical contributions that we believe will be of significant utility in the future study of intermediate degree, which, again, we view as a highly natural measure of plan complexity. One tool that we present (in Section 4) is a theorem showing how to convert an SPJ query plan to a relational structure along with a certain form of tree decomposition of the structure; this allows us to apply structural measures based on tree decompositions to reason about the intermediate degree of plans. While the way that our main theorem’s algorithm computes the minimized plan $p^{\prime}$ from a given plan $p$ is, on a high level, quite natural—essentially, $p^{\prime}$ is based on the core of the chase of a relational structure naturally derived from $p$—the proof that $p^{\prime}$ is minimal in the desired sense is nontrivial, since we have to reason about arbitrary plans that are semantically equivalent to $p^{\prime}$. The core is, intuitively speaking, a minimized version of a relational structure; the chase is a well-known procedure that enforces constraints on a relational structure.
Related work. Our work is inspired by and is closely related to the work of Atserias, Grohe, and Marx (Grohe and Marx, 2014; Atserias et al., 2013), as well as the follow-up work thereof by Gottlob, Lee, Valiant, and Valiant (Gottlob et al., 2012). A main result shown by (Atserias et al., 2013) yields a general upper bound—the AGM bound—on the output size of any pure join plan $q$, namely, a bound of $M_{\mathbf{D}}^{\rho^{*}(q)}$, where $\rho^{*}(q)$ is the so-called fractional edge cover number of the plan $q$. They also supply an argument that this upper bound is tight. Here, by a pure join plan, we mean a plan that consists only of joins and basic relations of a database.
Let us define the output degree of a query to be the minimum degree of a polynomial bounding the query’s output relation, as a function of the maximum relation size $M_{\mathbf{D}}$ of a database $\mathbf{D}$. Note that while the output degree is clearly invariant under passing from a query to another query that is semantically equivalent to the first, the intermediate degree does not enjoy this invariance. The AGM bound along with the argument that it is tight establish $\rho^{*}(q)$ as the output degree of a pure join plan $q$. The article (Atserias et al., 2013) also exhibits, for each such pure join plan, a semantically equivalent join-project plan whose intermediate relations all have size less than (or equal to) the output size, implying a characterization of the minimum intermediate degree (over semantically equivalent plans) as the fractional edge cover number—for pure join plans. We view the article (Atserias et al., 2013) as begging the study of more general classes of plans beyond pure join plans: they prove explicitly the necessity of using projections in the plans witnessing that the intermediate relation sizes can be made to be at most the output size!
Extending the AGM bound, the work (Gottlob et al., 2012) characterizes the output degree of any SPJ plan and considering databases where unary key constraints can be expressed; this is done by identifying a measure called the color number. Note that this measure generalizes the fractional edge cover number (as discussed in Section 4 of (Gottlob et al., 2012)). Our work builds on this measure and the accompanying understanding of unary key constraints. In particular, in the present work, we naturally extend the notion of color number to define a measure of tree decompositions. We remark that the results of (Gottlob et al., 2012) are phrased in terms of conjunctive queries, which are semantically equivalent to SPJ plans.
An important and influential line of research emerging after the work of Atserias, Grohe, and Marx presented new join algorithms, so-called worst-case optimal join algorithms, with running times shown to be bounded above by the AGM bound—apart from a multiplicative overhead (see for example (Ngo et al., 2013, 2018; Ngo, 2018) and the discussion therein). Building on this work, further research presented algorithms, with running time analyses, for more general forms of queries, including queries with aggregates (Joglekar et al., 2016; Khamis et al., 2016). The focus of these further works are on presenting algorithms and upper bounding their running time in terms of structural measures, generally without considerations of semantic equivalence; the present article establishes a complementary result by showing, for each query plan $p$ (of the specified form), a tight lower bound on the running time of a restricted class of algorithms, namely, algorithms that can be described by plans that are semantically equivalent to the given plan $p$.
2. Preliminaries
We use $\mathbb{Q}$ to denote the rational numbers, $\mathbb{Q}^{+}$ to denote the non-negative rational numbers, and $\mathbb{N}$ to denote the natural numbers, which we understand to include 00. For a natural number $k\in\mathbb{N}$, we use $[k]$ to denote the first $k$ positive natural numbers, that is, the set $\{i\in\mathbb{N}~{}|~{}1\leq i\leq k\}$. For a map $f:A\to B$ and some $S\subseteq A$, $f\upharpoonright S$ denotes the restriction of $f$ to $S$. For a tuple $\overline{a}=(a_{1},\ldots,a_{k})$ over a set $A$ and a map $f:A\to B$, by extension of notation, we write $f(\overline{a})$ for the entry-wise action of $f$ on the tuple, i.e., to denote $(f(a_{1}),\ldots,f(a_{k}))$; we further write $\{\overline{a}\}$ for the set $\{a_{1},\ldots,a_{k}\}$. When $A$ is a set, we use $A^{*}$ to denote the set of all finite-length tuples over $A$. When $\approx$ is a binary relation on a set $A$, we define $\approx^{*}$ as the reflexive-symmetric-transitive closure of $\approx$, that is, $\approx^{*}$ is the smallest equivalence relation on $A$ that contains $\approx$ as a subset.
2.1. Structures
A signature $\sigma$ is a finite set of relation symbols with associated arity. We denote by $R/n$ that symbol $R$ has arity $n\in\mathbb{N}$, we may also write $\mathrm{ar}(R)$ for $n$. A structure $\mathbf{A}$ over a signature $\sigma$ consists of a finite set called the universe or domain of the structure, denoted $\mathsf{dom}(\mathbf{A})$, and, for each symbol $R\in\sigma$, a relation $R^{\mathbf{A}}$ over $\mathsf{dom}(\mathbf{A})$ whose arity is that of $R$. An isolated element of a structure $\mathbf{A}$ is a universe element $a\in\mathsf{dom}(\mathbf{A})$ such that $a$ does not appear in any tuple of any relation of $\mathbf{A}$. An open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$ is a pair consisting of a structure $\mathbf{A}$ over signature $\sigma$ with no isolated elements and a tuple $\overline{a}$ over the universe $\mathsf{dom}(\mathbf{A})$ of $\mathbf{A}$. For a structure $\mathbf{A}$ over signature $\sigma$, let $M_{\mathbf{A}}$ be the maximum size over all relations of $\mathbf{A}$, that is, $\max_{R\in\sigma}|R^{\mathbf{A}}|$.
Two (open) structures are similar if they are defined over the same signature. When $\mathbf{A}$ and $\mathbf{B}$ are similar structures over the signature $\sigma$, we define their union, denoted by $\mathbf{A}\cup\mathbf{B}$, as the structure with universe $\mathsf{dom}(\mathbf{A})\cup\mathsf{dom}(\mathbf{B})$ and where, for each symbol $R\in\sigma$, it holds that $R^{\mathbf{A}\cup\mathbf{B}}=R^{\mathbf{A}}\cup R^{\mathbf{B}}$. For two similar structures $\mathbf{A}$ and $\mathbf{B}$ over a signature $\sigma$, a homomorphism from $\mathbf{A}$ to $\mathbf{B}$ is a map $h:\mathsf{dom}(\mathbf{A})\to\mathsf{dom}(\mathbf{B})$ such that for each symbol $R\in\sigma$, it holds that $\overline{a}\in R^{\mathbf{A}}$ implies $h(\overline{a})\in R^{\mathbf{B}}$. For similar open structures $(\mathbf{A},\overline{a})$ and $(\mathbf{B},\overline{b})$ with $\overline{a}$ and $\overline{b}$ having the same arity, a homomorphism $h$ from $(\mathbf{A},\overline{a})$ to $(\mathbf{B},\overline{b})$ is a homomorphism from $\mathbf{A}$ to $\mathbf{B}$ such that $h(\overline{a})=\overline{b}$. Two similar structures $\mathbf{A}$ and $\mathbf{B}$ are homomorphically equivalent if there exist a homomorphism from $\mathbf{A}$ to $\mathbf{B}$ and a homomorphism from $\mathbf{B}$ to $\mathbf{A}$. An isomorphism from $\mathbf{A}$ to $\mathbf{B}$ is a bijection $h:\mathsf{dom}(\mathbf{A})\to\mathsf{dom}(\mathbf{B})$ such that $h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{B}$, and $h^{-1}$ is a homomorphism from $\mathbf{B}$ to $\mathbf{A}$. Let $(\mathbf{A},\overline{a})$ be an open structure and let $\mathbf{D}$ be a structure that is similar to $\mathbf{A}$; we use $\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$ to denote the relation
$\{h(\overline{a})~{}|~{}\mbox{$h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$}\}.$
We conceive of this relation as the result of evaluating the open structure on the structure $\mathbf{D}$. Similarly, when $\mathbf{A}$ and $\mathbf{D}$ are similar structures and $S\subseteq\mathsf{dom}(\mathbf{A})$, we use $\mathsf{homs}(\mathbf{A},S,\mathbf{D})$ to denote the set of maps
$\{h\upharpoonright S~{}|~{}\mbox{$h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$}\}.$
Example 2.1.
As a small example, let $\sigma$ be the signature containing a single relation $E$ of arity $2$, and let $\mathbf{A}_{0}$ be the structure with universe $\mathsf{dom}(\mathbf{A}_{0})=\{u,v_{1},v_{2}\}$ and with relation $E^{\mathbf{A}_{0}}=\{(u,v_{1}),(u,v_{2})\}$. Consider the open structure $(\mathbf{A}_{0},(u))$; when $\mathbf{D}$ is a directed graph, we have that $\mathsf{homs}(\mathbf{A}_{0},(u),\mathbf{D})$ is equal to the arity $1$ relation containing each vertex (of $\mathbf{D}$) having an outgoing edge. Next, consider the open structure $(\mathbf{A}_{0},(v_{1},v_{2}))$; when $\mathbf{D}$ is a directed graph, we have that $\mathsf{homs}(\mathbf{A}_{0},(v_{1},v_{2}),\mathbf{D})$ is equal to the arity $2$ relation containing each pair of vertices (of $\mathbf{D}$) that each receive an incoming edge from a common vertex. This relation is symmetric, and when $x$ is a vertex of $\mathbf{D}$, this relation contains the pair $(x,x)$ if and only if $x$ receives an incoming edge (from some vertex).
The next fact follows from the classic work of Chandra and Merlin (Chandra and Merlin, 1977).
Proposition 2.2 ((Chandra and Merlin, 1977)).
Let $(\mathbf{A},\overline{a})$ and $(\mathbf{A}^{\prime},\overline{a^{\prime}})$ be open structures over some signature $\sigma$. There exists a homomorphism from $(\mathbf{A},\overline{a})$ to $(\mathbf{A}^{\prime},\overline{a^{\prime}})$ if and only if for each structure $\mathbf{D}$ over $\sigma$ we have that $\mathsf{homs}(\mathbf{A}^{\prime},\overline{a^{\prime}},\mathbf{D})\subseteq\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$.
For similar structures $\mathbf{A}$ and $\mathbf{B}$, we say that $\mathbf{A}$ is a substructure of $\mathbf{B}$ if $\mathsf{dom}(\mathbf{A})\subseteq\mathsf{dom}(\mathbf{B})$ and for each symbol $R\in\sigma$, it holds that $R^{\mathbf{A}}\subseteq R^{\mathbf{B}}$. When $\mathbf{B}$ is a structure and $S\subseteq\mathsf{dom}(\mathbf{B})$, the induced substructure $\mathbf{B}[S]$ is defined as the structure with universe $S$ and where, for each symbol $R$, it holds that $R^{\mathbf{B}[S]}=\{(b_{1},\ldots,b_{k})\in R^{\mathbf{B}}~{}|~{}\{b_{1},\ldots,b_{k}\}\subseteq S\}$. When $\mathbf{A}$ is a substructure of $\mathbf{B}$, a retraction from $\mathbf{B}$ to $\mathbf{A}$ is a homomorphism from $\mathbf{B}$ to $\mathbf{A}$ that acts as the identity on $\mathsf{dom}(\mathbf{A})$. When there exists a retraction from $\mathbf{B}$ to $\mathbf{A}$, we say that $\mathbf{B}$ retracts to $\mathbf{A}$. A core of a structure $\mathbf{B}$ is a substructure $\mathbf{C}$ of $\mathbf{B}$ such that $\mathbf{B}$ retracts to $\mathbf{C}$, but $\mathbf{C}$ does not retract to any proper substructure of $\mathbf{C}$, that is, any substructure of $\mathbf{C}$ whose universe is a proper subset of $\mathbf{C}$’s universe. A structure is a core if it is a core of itself.
Example 2.3.
Consider the structure $\mathbf{A}_{0}$ of Example 2.1. For each $i=1,2$, let $\mathbf{A}_{i}$ be the structure with universe $\mathsf{dom}(\mathbf{A}_{i})=\{u,v_{i}\}$ and with relation $E^{\mathbf{A}_{i}}=\{(u,v_{i})\}$. Clearly, each of $\mathbf{A}_{1},\mathbf{A}_{2}$ is a substructure of $\mathbf{A}_{0}$. In addition, we have that the map fixing $u$ and $v_{1}$ and sending $v_{2}$ to $v_{1}$ is a retraction from $\mathbf{A}_{0}$ to $\mathbf{A}_{1}$; dually, the map fixing $u$ and $v_{2}$ and sending $v_{1}$ to $v_{2}$ is a retraction from $\mathbf{A}_{0}$ to $\mathbf{A}_{2}$. Each of $\mathbf{A}_{1},\mathbf{A}_{2}$ is a core of $\mathbf{A}_{0}$ (and of itself): to argue this, we need to argue that (say) $\mathbf{A}_{1}$ has no retraction to a proper substructure $\mathbf{A}^{\prime}$ (of $\mathbf{A}_{1}$). (With respect to $\mathbf{A}_{0}$, the structures $\mathbf{A}_{1}$, $\mathbf{A}_{2}$ are symmetric to each other.) We argue this as follows: if we had such a proper substructure, it would have a size $1$ universe, and the retraction would map each element in $\mathsf{dom}(\mathbf{A}_{1})$ to the single element $x$ in that size $1$ universe. But then this proper substructure $\mathbf{A}^{\prime}$ would have $E^{\mathbf{A}^{\prime}}=\{(x,x)\}$, which could not be a subset of $E^{\mathbf{A}_{1}}$, contradicting that $\mathbf{A}^{\prime}$ is a substructure of $\mathbf{A}_{1}$.
The following facts are known and straightforward to verify.
Proposition 2.4.
The following facts hold.
  1. (1)
    For each structure $\mathbf{B}$, there exists a core of $\mathbf{B}$.
  2. (2)
    A core $\mathbf{C}$ of a structure $\mathbf{B}$ is homomorphically equivalent to $\mathbf{B}$.
  3. (3)
    Any two cores of a structure are isomorphic. (We thus, by a mild abuse of terminology, speak of the core of a structure.)
  4. (4)
    When $\mathbf{C}$ is a core, and $\mathbf{B}$ is a structure that is homomorphically equivalent to $\mathbf{C}$, there exists a core $\mathbf{C}^{\prime}$ of $\mathbf{B}$ that is isomorphic to $\mathbf{C}$.
An open structure is not a structure per se. However, it will be useful to translate an open structure to a structure. We reserve a set $\{R_{k}\}_{k\in\mathbb{N}}$ of relation names for the purpose of translating open structures to structures, where each relation $R_{k}$ has arity $k$ and is distinct from the relations mentioned in any non-augmented signature. When $(\mathbf{A},(a_{1},\ldots,a_{k}))$ is an open structure over the signature $\sigma$, we define its augmented structure, denoted by $\mathsf{aug}(\mathbf{A},(a_{1},\ldots,a_{k}))$, as the structure $\mathbf{A}_{+}$ over signature $\sigma\cup\{R_{k}\}$ where $R^{\mathbf{A}_{+}}=R^{\mathbf{A}}$ for each $R\in\sigma$, and $R_{k}^{\mathbf{A}_{+}}=\{(a_{1},\ldots,a_{k})\}$. In particular, this means that if open structures $(\mathbf{A},\overline{a})$ and $(\mathbf{A}^{\prime},\overline{a^{\prime}})$ are similar and tuples $\overline{a}$ and $\overline{a^{\prime}}$ have the same arity, then also the augmented structures $\mathsf{aug}(\mathbf{A},\overline{a})$ and $\mathsf{aug}(\mathbf{A}^{\prime},\overline{a^{\prime}})$ are similar.
We associate each open structure with its augmented structure, so for example we regard an open structure $(\mathbf{C},\overline{c})$ as a core of an open structure $(\mathbf{B},\overline{b})$ when $\mathsf{aug}(\mathbf{C},\overline{c})$ is a core of $\mathsf{aug}(\mathbf{B},\overline{b})$. Note that a homomorphism from an open structure $(\mathbf{A},\overline{a})$ to an open structure $(\mathbf{B},\overline{b})$, as we defined it above, is a homomorphism from $\mathsf{aug}(\mathbf{A},\overline{a})$ to $\mathsf{aug}(\mathbf{B},\overline{b})$, and vice-versa.
2.2. Query plans
For completeness and clarity, we precisely define the types of select-project-join plans (SPJ plans) to be studied. For each $m\geq 0$, we define an $m$-suitable identification as an expression $j=k$ where $j,k\in[m]$.
For a signature $\sigma$, a SPJ plan over $\sigma$ is inductively defined as follows:
  • (basic) For each $R\in\sigma$, it holds that $R$ is a plan of arity $\mathrm{ar}(R)$.
  • (select) When $p$ is a plan of arity $m$ and $\theta$ is a set of $m$-suitable identifications, $\sigma_{\theta}(p)$ is a plan of arity $m$.
  • (project) When $p$ is a plan of arity $m$ and $j_{1},\ldots,j_{n}$ is a sequence of numbers from $[m]$, with $n\geq 0$, it holds that $\pi_{j_{1},\ldots,j_{n}}(p)$ is a plan of arity $n$.
  • (join) When $p_{1},\ldots,p_{\ell}$ are plans of arity $m_{1},\ldots,m_{\ell}$, respectively, with $\ell\geq 1$, and $\theta$ is a set of $(m_{1}+\cdots+m_{\ell})$-suitable identifications, $\Join_{\theta}(p_{1},\ldots,p_{\ell})$ is a plan of arity $(m_{1}+\cdots+m_{\ell})$.
Suppose that $Q$ is a relation of arity $m$. For any set $\theta$ of $m$-suitable identifications, we define
$\sigma_{\theta}(Q)=\{(d_{1},\ldots,d_{m})\in Q~{}|~{}\mbox{for each $(j=k)\in\theta$, $d_{j}=d_{k}$ holds}\}.$
For any sequence $j_{1},\ldots,j_{n}\in[m]$, we define
$\pi_{j_{1},\ldots,j_{n}}(Q)=\{(d_{j_{1}},\ldots,d_{j_{n}})~{}|~{}(d_{1},\ldots,d_{m})\in Q\}.$
Let $p$ be a SPJ plan over a signature $\sigma$, let $\mathbf{D}$ be a structure over $\sigma$, and let $m$ be the arity of $p$. We define the output of $p$ over $\mathbf{D}$, denoted by $\mathsf{out}(p,\mathbf{D})$, to be the arity $m$ relation over $D$ defined as follows. Note that, as is usual in this context, we understand that there is a unique tuple of arity 00, called the empty tuple.
  • When $p$ has the form $R$, we define $\mathsf{out}(p,\mathbf{D})=R^{\mathbf{D}}.$
  • When $p$ has the form $\sigma_{\theta}(p^{\prime})$, we define $\mathsf{out}(p,\mathbf{D})=\sigma_{\theta}(\mathsf{out}(p^{\prime},\mathbf{D})).$
  • When $p$ has the form $\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, we define $\mathsf{out}(p,\mathbf{D})=\pi_{j_{1},\ldots,j_{n}}(\mathsf{out}(p^{\prime},\mathbf{D})).$
  • When $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, we define $\mathsf{out}(p,\mathbf{D})=\sigma_{\theta}(\mathsf{out}(p_{1},\mathbf{D})\times\cdots\times\mathsf{out}(p_{\ell},\mathbf{D})).$
Observe that selection is a special case of join, in particular, it corresponds to a $1$-way join. We say that two plans $p,p^{\prime}$ over the same signature $\sigma$ are semantically equivalent when, for every structure $\mathbf{D}$ over $\sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{\prime},\mathbf{D})$. Further, for a set $\Sigma$ of keys over $\sigma$, we say that two plans $p,p^{\prime}$ over $\sigma$ are $\Sigma$-semantically equivalent when, for every structure $\mathbf{D}$ over $\sigma$ that satisfies $\Sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{\prime},\mathbf{D})$. A subplan of a plan is defined in the usual manner. Formally, we define a set $\mathsf{subplans}(p)$, for each plan $p$, inductively as follows:
  • When $p$ has the form $R$, we define $\mathsf{subplans}(p)=\{p\}$.
  • When $p$ has the form $\sigma_{\theta}(p^{\prime})$ or $\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, we define $\mathsf{subplans}(p)=\mathsf{subplans}(p^{\prime})\cup\{p\}.$
  • When $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, we define $\mathsf{subplans}(p)=(\bigcup_{i=1}^{\ell}\mathsf{subplans}(p_{i}))\cup\{p\}.$
We then say that $q$ is a subplan of $p$ when $q\in\mathsf{subplans}(p)$. Towards establishing our positive results, that is, presenting SPJ plans satisfying certain upper bounds, we rely on the notion of well-behaved plans. A plan $p$ is well-behaved if each subplan $q$ that is a join $\Join_{\theta}(p_{1},\ldots,p_{\ell})$ of plans $p_{1},\ldots,p_{\ell}$ having arities $m_{1},\ldots,m_{\ell}$, letting $s$ denote the sum $m_{1}+\cdots+m_{\ell}$, the following is satisfied: there exists an index $k\in[s]$ such that for each index $i\in[s]\setminus([m_{1}]\cup\{k\})$, there exists an index $j\in[m_{1}]\cup\{k\}$ such that $(i=j)\in\theta$. In words, each join adds at most one column of new information to the plan $p_{1}$: letting $J$ denote the set containing (1) the indices in $[m_{1}]$, which correspond to $p_{1}$, and (2) another chosen index $k$; it must be that every index $i\notin J$ is equated with an index in $J$. This property allows for the efficient computation of such joins and, thus, well-behaved plans; we will make this formal in Proposition 3.4.
2.3. Hypergraphs and tree decompositions
A hypergraph is a pair $(V,E)$ where $V$ is a set, called the vertex set, and $E\subseteq 2^{V}$ is a set called the edge set. A graph is a hypergraph where each edge has size $2$. When $H$ is a hypergraph, we use $V(H)$ and $E(H)$ to denote its vertex set and edge set, respectively. An isolated vertex of a hypergraph $H$ is a vertex not occurring in any edge, that is, an element of $V(H)\setminus\bigcup_{e\in E(H)}e$. When $H$ is a hypergraph and $S\subseteq V(H)$, we use $H[S]$ to denote the induced subhypergraph of $H$ on $S$, that is, the hypergraph with vertex set $S$ and with edge set $\{e\cap S~{}|~{}e\in E(H)\}$.
When $\mathbf{A}$ is a structure over signature $\sigma$ we use $\mathbf{H}(\mathbf{A})$ to denote the hypergraph with vertex set $\mathsf{dom}(\mathbf{A})$ and edge set $\{\{\overline{a}\}~{}|~{}R\in\sigma,\overline{a}\in R^{\mathbf{A}}\}$. When $(\mathbf{A},\overline{a})$ is an open structure, we use $\mathbf{H}(\mathbf{A},\overline{a})$ to denote the hypergraph with vertex set $\mathsf{dom}(\mathbf{A})$ and with edge set $E(\mathbf{H}(\mathbf{A}))\cup\{\{\overline{a}\}\}$.
A tree decomposition of a hypergraph $H$ is a pair $(T,\chi)$, where $T$ is a tree and $\chi$ is a labeling function $V(T)\rightarrow 2^{V(H)}$, i.e., it assigns each vertex of $T$ a subset of vertices of $H$, such that
  1. (1)
    (vertex coverage) for each $v\in V(H)$, there exists $t\in V(T)$ such that $v\in\chi(t)$,
  2. (2)
    (edge coverage) for each $e\in E(H)$, there exists $t\in V(T)$ such that $e\subseteq\chi(t)$, and
  3. (3)
    (connectivity) for each $v\in V(H)$, the set $\{t\in V(T)~{}|~{}v\in\chi(t)\}$ induces a connected subtree of $T$.
Note that, if $H$ has no isolated vertices, condition $(1)$ is implied by condition $(2)$. A tree decomposition of an open structure $(\mathbf{A},\overline{a})$, is a tree decomposition of its hypergraph $\mathbf{H}(\mathbf{A},\overline{a})$.
2.4. Key constraints
A key constraint $\kappa$ over signature $\sigma$ is an expression $\textrm{key}(R)=K$, where $R/n\in\sigma$ and $K\subseteq[n]$. For an $n$-tuple $\overline{a}=(a_{1},\ldots,a_{n})$ and some $K=\{i_{1},\ldots,i_{m}\}\subseteq[n]$, for some $m\in[n]$, we write $\overline{a}[K]$ for the tuple $(a_{i_{1}},\ldots,a_{i_{m}})$. A structure $\mathbf{A}$ over signature $\sigma$ satisfies $\kappa$ if, for every two tuples $\overline{a},\overline{b}\in R^{\mathbf{A}}$, $\overline{a}[K]=\overline{b}[K]$ implies $\overline{a}=\overline{b}$. If, for some key $\kappa$ of the form $\textrm{key}(R)=K$, the set $K$ is a singleton, we say that $\kappa$ is a unary key. We say a structure satisfies a set of keys $\Sigma$, if it satisfies every key $\kappa\in\Sigma$.
The chase is a well-known tool for enforcing key dependencies (see, e.g., (Maier et al., 1979; Deutsch et al., 2006)). For an open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$ and a set $\Sigma$ of keys over $\sigma$ a chase step on $(\mathbf{A},\overline{a})$ w.r.t. $\Sigma$ is a substitution of the form $x\mapsto y$, which is applied globally on $\mathbf{A}$ and $\overline{a}$, where there are two tuples $\overline{b},\overline{b^{\prime}}\in R^{\mathbf{A}}$ and a key $\textrm{key}(R)=K$ for some $R/n\in\sigma$, such that $\overline{b}[K]=\overline{b^{\prime}}[K]$ but there exists some $i\in[n]\setminus K$ such that $\overline{b}[i]=x$ and $\overline{b^{\prime}}[i]=y$. The chase of $(\mathbf{A},\overline{a})$ w.r.t. $\Sigma$, denoted $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, is the result of repeated applications of chase steps until a fixpoint is reached; this fixpoint is known to be unique up to isomorphism. The chase of a structure $\mathbf{A}$ over signature $\sigma$ w.r.t. $\Sigma$, denoted $\textrm{Chase}_{\Sigma}(\mathbf{A})$, is defined analogously in the obvious way.
We next define a measure for hypergraphs based on a notion, called color number, defined on conjunctive queries and introduced in (Gottlob et al., 2012). For a structure $\mathbf{A}$ over signature $\sigma$ and a set $\Sigma$ of keys over $\sigma$, a valid coloring of $\mathbf{A}$ w.r.t. $\Sigma$ is a mapping $\mathsf{Col}_{\Sigma}^{\mathbf{A}}:\mathsf{dom}(\mathbf{A})\rightarrow 2^{\{1,\ldots,k\}}$, for some $k>0$, which assigns to each element of the universe of $\mathbf{A}$ a subset of $k$ colors such that the following hold:
  1. (1)
    For each key $\textrm{key}(R)=\{i_{1},\ldots,i_{m}\}\in\Sigma$, i.e., $R/n\in\sigma$ and $\{i_{1},\ldots,i_{m}\}\subseteq[n]$, we have that for every $\overline{a}\in R^{\mathbf{A}}$ and every $i\in[n]\setminus\{i_{1},\ldots,i_{m}\}$ it holds that $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(\overline{a}[i])\subseteq\bigcup_{j\in[m]}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(\overline{a}[i_{j}])$.
  2. (2)
    There exists an element $x\in\mathsf{dom}(\mathbf{A})$ such that $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)\neq\emptyset$.
For a structure $\mathbf{A}$ over signature $\sigma$, a set $S\subseteq\mathsf{dom}(\mathbf{A})$, and a set $\Sigma$ of keys over $\sigma$ the color number of $S$ w.r.t. $\mathbf{A}$ and $\Sigma$, denoted $C_{\Sigma}^{\mathbf{A}}(S)$, is defined as the maximal ratio of colors assigned to the elements in $S$ over the maximum number of colors assigned to the elements of some tuple in $\mathbf{A}$ which is attainable by any valid coloring (for any $k$) of $\mathbf{A}$ w.r.t. $\Sigma$. Formally, where $\max_{\mathsf{Col}_{\Sigma}^{\mathbf{A}}}$ ranges over all such valid colorings:
$C_{\Sigma}^{\mathbf{A}}(S)=\max_{\mathsf{Col}_{\Sigma}^{\mathbf{A}}}\frac{|\bigcup_{x\in S}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)|}{\max_{\overline{b}\in\mathbf{A}}|\bigcup_{x\in\overline{b}}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)|}.$
A crucial observation is that for a coloring that maximises the color number of $S$, we have that $x\notin S$ implies $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)=\emptyset$. Note that, for a structure $\mathbf{A}$ over signature $\sigma$, the color number $C_{\Sigma}^{\mathbf{A}}(S)$, as defined here, is equal to the color number, as defined in (Gottlob et al., 2012), of the conjunctive query whose atoms are those in $\{R(v_{1},\ldots,v_{k})~{}|~{}R\in\sigma,(v_{1},\ldots,v_{k})\in R\}$, and whose output variables are $S$. We refer the reader to (Gottlob et al., 2012) for more information on the color number.
Let us view $C_{\Sigma}^{\mathbf{A}}$ as a mapping from the power set of $\mathsf{dom}(\mathbf{A})$ to $\mathbb{Q}^{+}$. We define the following width notions. For an open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$, a set $\Sigma$ of keys over $\sigma$, and a tree decomposition $F=(T,\chi)$ of $\mathbf{H}(\mathbf{A},\overline{a})$, let $C_{\Sigma}^{\mathbf{A}}\textup{-width}(F)=\max\{C_{\Sigma}^{\mathbf{A}}(\chi(t))~{}|~{}t\in V(T)\}$. Further, let $C_{\Sigma}^{\mathbf{A}}\textup{-width}(\mathbf{A},\overline{a})=\min_{F}C_{\Sigma}^{\mathbf{A}}\textup{-width}(F)$, where $F$ ranges over all tree decompositions of $\mathbf{H}(\mathbf{A},\overline{a})$.
3. Intermediate Degree and Main Theorem Statement
A key measure in the context of query optimization is the size of the output of a query. Here, we are interested in an asymptotic measure: the output degree of a query plan $p$, which, informally, is the smallest possible degree of a polynomial bounding the output size of $p$ on any structure $\mathbf{D}$ satisfying given constraints, as a function of the maximum relation size $M_{\mathbf{D}}$; the formal definition is the following.
Definition 3.1 (Output Degree).
Let $\Sigma$ be a set of keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$, and let $d\in\mathbb{Q}$. The plan $p$ has output degree $\leq d$ if there exists a function $h:\mathbb{N}\to\mathbb{N}$, with $h(M)\in O(M^{d})$, such that for every structure $\mathbf{D}$ satisfying $\Sigma$, it holds that $|\mathsf{out}(p,\mathbf{D})|\leq h(M_{\mathbf{D}})$. We say that $p$ has output degree $d$ if it has output degree $\leq d$, but for all $\epsilon>0$, it does not have output degree $\leq(d-\epsilon)$.
Instead of just focusing on some plan as a whole, we are interested in bounding also the size of the output of every subplan of the original plan, or in other words, bounding the size of intermediate results obtained during the execution of the plan. To do so, we say a plan $p$ has intermediate degree $d$ if there exists a degree $d$ polynomial bounding the sizes of all intermediate relations computed during the execution of $p$ on any structure $\mathbf{D}$ satisfying given constraints, as a function of $M_{\mathbf{D}}$—and again, there is no lower degree polynomial providing such a bound.
Definition 3.2 (Intermediate Degree).
Let $\Sigma$ be a set of keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$, and let $d\in\mathbb{Q}$. The plan $p$ has intermediate degree $\leq d$ if there exists a function $h:\mathbb{N}\to\mathbb{N}$, with $h(M)\in O(M^{d})$, such that for every structure $\mathbf{D}$ satisfying $\Sigma$, and for every subplan $q$ of $p$, it holds that $|\mathsf{out}(q,\mathbf{D})|\leq h(M_{\mathbf{D}})$. We say that $p$ has intermediate degree $d$ if it has intermediate degree $\leq d$, but for all $\epsilon>0$, it does not have intermediate degree $\leq(d-\epsilon)$.
Note that the above definitions of output degree and intermediate degree are with respect to the keys $\Sigma$, but we omit $\Sigma$ when it is clear from context for the sake of readability. Given a set of constraints $\Sigma$, we say that an SPJ plan $p$ has the best possible intermediate degree if there exists some $d\in\mathbb{Q}$ such that $p$ has intermediate degree $d$, and for every SPJ plan $p^{\prime}$ that is $\Sigma$-semantically equivalent to $p$, and there is no value $\epsilon>0$ such that the plan $p^{\prime}$ has intermediate degree $\leq(d-\epsilon)$. In other words, no $\Sigma$-semantically equivalent plan achieves an intermediate degree that is strictly lower than that of $p$.
We analyze a plan $p$ by means of its so called $p$-representation, an open structure $(\mathbf{A},\overline{a})$ for which it holds that, for any structure $\mathbf{D}$, we have that $\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$. A formal treatment of $p$-representation and related notions utilized to show our results will follow in Section 4. We are now ready to formulate our main theorem.
Theorem 3.3 (Main Theorem).
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ be a $p$-representation. Let $(\mathbf{C},\overline{c})$ be a core of $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$; then there exists a well-behaved plan $p^{\prime}$ where
  • $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$,
  • $p^{\prime}$ has intermediate degree $d$, and
  • $p^{\prime}$ has the best possible intermediate degree.
There exists an exponential-time algorithm that, given an SPJ plan $p$, outputs such a plan $p^{\prime}$; the plan $p^{\prime}$ has size polynomial in the size of $p$.
We remark that given an SPJ plan $p$ the value $d$ is computable in exponential time; this follows quite directly from the polynomial time computability of the color number, which follows from results in Sections 3 and 4 from (Gottlob et al., 2012). The algorithm provided by Theorem 3.3 passes from a plan $p$ to a second plan $p^{\prime}$ with the best possible intermediate degree. A natural question that arises is how much time it takes to evaluate $p^{\prime}$ on a given structure $\mathbf{D}$. The following proposition, which can be argued elementarily, gives an answer to this question: given a well-behaved plan, apart from a multiplicative overhead depending on the plan, the time needed is at most $|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d}$, where $d$ is the intermediate degree of the plan—so in essence, the intermediate degree of the plan dictates the polynomial degree of evaluation.
Proposition 3.4.
There exists an algorithm that, given a well-behaved SPJ plan $p$ of intermediate degree $\leq d$ and a structure $\mathbf{D}$, evaluates $p$ on $\mathbf{D}$ within time $O(|p|^{2}*|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d})$.
To argue this proposition, we consider an algorithm that evaluates the plan $p$ inductively. It suffices to show that each subplan $q\in\mathsf{subplans}(p)$ can be computed in time $O(|p|*|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d})$. By the definition of intermediate degree, each subplan $q$ has at most $O(|M_{\mathbf{D}}|^{d})$ answers (tuples in $\mathsf{out}(p,\mathbf{D})$). When $q$ is a basic plan, the required bound holds immediately. When $q$ is a select or project subplan, i.e., of the form $\sigma_{\theta}(q^{\prime})$ or $\pi_{j_{1},\ldots,j_{n}}(q^{\prime})$, respectively, the answers of $q$ can be naturally derived from the answers of the immediate subplan $q^{\prime}$, within the given time bound. The key point is that, when $q$ is a join subplan, by well-behavedness, it is of the form $\Join_{\theta}(q_{1},\ldots,q_{\ell})$ and $\mathsf{out}(q,\mathbf{D})$ can be computed by maintaining a relation that is a subset of $\mathsf{out}(q_{1},\mathbf{D})\times\mathsf{dom}(\mathbf{D})$. When evaluating the multiway join, we can thus evaluate the sequence of binary joins $\Join_{\theta_{\ell-1}}(\ldots(\Join_{\theta_{2}}(\Join_{\theta_{1}}(q_{1},q_{2}),q_{3}),\ldots),q_{\ell})$, where $\theta_{1},\ldots,\theta_{\ell-1}$ are the corresponding equalities from $\theta$ for the join partners in question. Each binary join can be evaluated in time $O(|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d})$ by (Flum et al., 2002), and in total, there are at most $|p|$ many joins, giving us the desired runtime of $O(|p|*|\mathsf{dom}(\mathbf{D})|*|M_{\mathbf{D}}|^{d})$ to evaluate the subplan $q$.
Note that our main theorem only considers constraints which are unary keys. This is a limitation that arises from the literature regarding the bounds on the size of the output of a plan and is challenging to overcome. To extend the bounds we inherit from the literature, one is likely to make significant progress towards a long-standing open problem in information theory (Gottlob et al., 2012).
Proving Theorem 3.3 essentially requires us to show two statements: (i) we can synthesize a plan, which is $\Sigma$-equivalent to the original one and has intermediate degree $\leq d$ and (ii) any plan which is $\Sigma$-equivalent to the original one has a subplan with output degree at least $d$. We will formalize and prove (i) and (ii) in Theorems 5.1 and 5.2, respectively, in Section 5.
To show the theorem, we make two more basic observations on the introduced degree notions. We begin by presenting a definition that will aid us in arguing that a plan has a certain output degree. We say that $p$ has output degree $\geq d$ if there exists a function $g:\mathbb{N}\to\mathbb{N}$ with $g(M)\in\Omega(M^{d})$ and there exists an infinite sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures with $M_{\mathbf{D}_{n}}$ unbounded such that (for all $n\geq 1$) $|\mathsf{out}(p,\mathbf{D}_{n})|\geq g(M_{\mathbf{D}_{n}})$. We then have the following lemma.
Lemma 3.5.
Let $d\geq 0$. If a plan has output degree $\geq d$, then for any $\epsilon>0$, it does not have output degree $\leq(d-\epsilon)$.
Next, we characterize the intermediate degree of a plan in terms of the output degrees of subplans, in the following sense.
Lemma 3.6.
Let $d\geq 0$. A plan $p$ has intermediate degree $\leq d$ if and only if each subplan $q$ of $p$ has output degree $\leq d$.
Theorem 3.3 can then be argued as follows: Given a plan $p$, we first compute the open structure $(\mathbf{A},\overline{a})$, which can be done in polynomial time by Theorem 4.4. Then, the algorithm of Theorem 3.3 computes $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, again in polynomial time. It then computes the core of this open structure; this can be performed in exponential time. Let $(\mathbf{C},\overline{c})$ be this core. Note that the size of $(\mathbf{C},\overline{c})$ is bounded above by the size of $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, since the core of a structure is a substructure of the structure. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. From $(\mathbf{C},\overline{c})$, we can then synthesize the desired plan $p^{\prime}$ of intermediate degree $\leq d$ by statement (i), formalized in Theorem 5.1, from above. Consider any plan $p^{\prime\prime}$ that is $\Sigma$-semantically equivalent to $p^{\prime}$. By statement (ii), formalized in Theorem 5.2, $p^{\prime\prime}$ has a subplan with output degree $\geq d$. For any $\epsilon>0$, by Lemma 3.5, this subplan does not have output degree $\leq(d-\epsilon)$, and so by Lemma 3.6, $p^{\prime\prime}$ does not have intermediate degree $\leq(d-\epsilon)$. Hence, we have computed a $\Sigma$-equivalent plan $p^{\prime}$ of polynomial size and of intermediate degree $d$, which is the best possible intermediate degree.
In the next section, we will introduce the connection from plans to open structures and auxiliary notions, which are needed to prove the theorems formalizing statements (i) and (ii) in Section 5.
4. Relating queries to structures
There is a known correspondence between Boolean SPJ plans and structures (Chandra and Merlin, 1977); by a Boolean SPJ plan, we mean one that evaluates to true or false (or alternatively, a relation of arity 00) on a given structure $\mathbf{D}$. This correspondence yields that each such plan can be translated to a structure $\mathbf{A}$ such that, for an arbitrary structure $\mathbf{D}$, the plan is true on $\mathbf{D}$ if and only if there is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$. We prove a version of this correspondence in Theorem 4.4 for SPJ plans that can output relations of any finite arity, and not just relations of arity 00. In order to formulate our version, we utilize the notion of open structure.
Definition 4.1 ($p$-Representation).
Let $p$ be an SPJ plan over signature $\sigma$. A $p$-representation is an open structure $(\mathbf{A},\overline{a})$ over $\sigma$, where for each structure $\mathbf{D}$ over $\sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$.
Apart from showing that any SPJ plan $p$ can be translated to a $p$-representation, we also want to provide a tree decomposition of this open structure with certain additional properties, which we call $p$-decomposition. This allows us to relate plans to structural, decomposition-based measures. While the translation that we present is, on a high level, along the lines of known translations from SPJ plans to conjunctive queries, we emphasize that our translation also provides a tree decomposition that will allow us to analyze the intermediate degree. In particular, for a given plan, each subplan will correspond to a vertex of the tree decomposition whose associated set of elements corresponds to the free variables of the subplan.
We say a tree decomposition $(T,\chi)$ is rooted if there is a distinguished node $v_{0}\in V(T)$, called the root of $T$. The formal definition of $p$-decomposition follows.
Definition 4.2 ($p$-Decomposition).
Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ a $p$-representation. A $p$-decomposition of $(\mathbf{A},\overline{a})$ is a rooted tree decomposition $E=(T,\chi)$ of $(\mathbf{A},\overline{a})$ such that there exists
  • a bijection $\alpha:\mathsf{subplans}(p)\to V(T)$, where $\alpha(p)$ is the root of $T$, and
  • a map $\beta:\mathsf{subplans}(p)\to\mathsf{dom}(\mathbf{A})^{*}$, where $\beta(p)$ is equal to $\overline{a}$,
where for each $q\in\mathsf{subplans}(p)$, we have that $\{\beta(q)\}=\chi(\alpha(q))$.
In the definition of $p$-decomposition, each subplan $q$ of $p$ corresponds to a vertex of the tree $T$, via $\alpha$, and for each such subplan $q$, it holds that $\beta(q)$ is a tuple that lists, in some order (and possibly with repetition), the elements of the set to which $\chi$ maps q’s tree vertex.
Let $p$ be an SPJ plan over signature $\sigma$, $(\mathbf{A},\overline{a})$ be a $p$-representation, and $E=(T,\chi)$ be a rooted $p$-decomposition $E=(T,\chi)$ of $(\mathbf{A},\overline{a})$ with associated mappings $\alpha$ and $\beta$. For a vertex $v\in T$, let $S_{v}$ be the union of $\chi(u)$ over all descendants $u$ of $v$ (including $v$ itself) and let $\mathbf{A}_{\leq v,E}$ denote the induced substructure $\mathbf{A}[S_{v}]$. We say a subplan $q\in\mathsf{subplans}(p)$ satisfies the containment property w.r.t. a $p$-decomposition if, for each structure $\mathbf{D}$ over $\sigma$, the relationship $\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}_{\leq\alpha(q),E},\beta(q),\mathbf{D})$ holds. Further, we say a $p$-decomposition satisfies the containment property, if every subplan $q\in\mathsf{subplans}(p)$ satisfies the containment property w.r.t. the $p$-decomposition. The containment property will allow us to lower bound the sizes of intermediate relations of $p$ using a $p$-representation along with a $p$-decomposition satisfying the containment property.
Example 4.3.
Consider the signature $\sigma$ containing the binary relations $R/2$ and $S/2$ and the SPJ plan $p=\Join_{(1=3)}(R,\pi_{1}(S))$ over $\sigma$. Then it can be readily seen, with structure $\mathbf{A}=\{R(a,b),S(a,d)\}$ and tuple $\overline{a}=(a,b,a)$, that $(\mathbf{A},\overline{a})$ is a $p$-representation. Consider the following trees:
On the left, we have the plan $p$ represented as a tree (each node represents the subplan of its subtree), while on the right we have some tree $T$. We provide mappings $\alpha$, $\beta$, and $\chi$ such that $(T,\chi)$ is a $p$-decomposition via mappings $\alpha$ and $\beta$. First, $\alpha$ is the obvious bijection from the subplans of $p$ to $V(T)$ with $\alpha(p)=v_{0}$, $\alpha(R)=v_{1}$, $\alpha(\pi_{1}(S))=v_{2}$, and $\alpha(S)=v_{3}$. Further, we define $\beta$ as $\beta(p)=\overline{a}$, $\beta(R)=(a,b)$, $\beta(\pi_{1}(S))=(a)$, and $\beta(S)=(a,d)$. Lastly, define $\chi(\alpha(q))$ as $\{\beta(q)\}$ for all subplans $q$ of $p$. It can be readily verified that $(T,\chi)$ is indeed a $p$-decomposition, and that it satisfies the containment property.
We show with the following theorem, that a $p$-representation and a $p$-decomposition with the above properties always exist and are efficiently computable.
Theorem 4.4.
There exists a polynomial-time algorithm that, given an SPJ plan $p$ over signature $\sigma$, outputs a $p$-representation over $\sigma$ along with a $p$-decomposition of this $p$-representation which satisfies the containment property.
[…]
Proof.
We show, by induction on the structure of $p$, how to derive an extended $p$-representation $(\mathbf{A},\overline{a},\approx)$ and $p$-decomposition thereof which satisfies the containment property. We then have that $(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle)$ is a $p$-representation along with a $p$-decomposition. We prove the statement inductively over the structure of $p$. Here, we show the case where $p$’s outer-most operator is a join; the other cases are relatively straightforward and treated in the appendix.
Suppose that $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, where each $p_{i}$ has arity $m_{i}$, for $i\in[\ell]$, and $p$ has arity $m=m_{1}+\cdots+m_{\ell}$. By induction, there exists (for each $i$) an extended $p_{i}$-representation $(\mathbf{A}_{i},\overline{t_{i}},\approx_{i})$. We may assume without loss of generality that the universes $\mathsf{dom}(\mathbf{A}_{i})$ are pairwise disjoint (if they are not, their elements can be renamed to achieve this). We let $\overline{a}=(a_{1},\ldots,a_{m})$ denote the tuple $\overline{t_{1}}\ldots\overline{t_{\ell}}$, that is, the tuple obtained by concatenating the entries of the tuples $\overline{t_{i}}$, $i\in[\ell]$, in order. Define $\approx_{0}$ as the relation $\bigcup_{i=1}^{\ell}\approx_{i}$, i.e., collecting all previously known equivalences, and define $\approx$ as the relation $\approx_{0}\cup\bigcup_{(j=k)\in\theta}\{(a_{j},a_{k})\}$, i.e., adding the equivalences newly introduced by $\theta$. We define $\mathbf{A}$ as $\mathbf{A}_{1}\cup\cdots\cup\mathbf{A}_{\ell}$. Then we have that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation, since for any structure $\mathbf{D}$ over $\sigma$:
$\mathsf{homs}(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle,\mathbf{D})$
$=\sigma_{\theta}(\mathsf{homs}(\mathbf{A}\langle\approx_{0}\rangle,\overline{a}\langle\approx_{0}\rangle,\mathbf{D}))$
$=\sigma_{\theta}(\mathsf{homs}(\mathbf{A}_{1}\langle\approx_{1}\rangle,\overline{t_{1}}\langle\approx_{1}\rangle,\mathbf{D})\times\cdots\times\mathsf{homs}(\mathbf{A}_{\ell}\langle\approx_{\ell}\rangle,\overline{t_{\ell}}\langle\approx_{\ell}\rangle,\mathbf{D}))$
$=\sigma_{\theta}(\mathsf{out}(p_{1},\mathbf{D})\times\cdots\times\mathsf{out}(p_{\ell},\mathbf{D}))$
$=\mathsf{out}(p,\mathbf{D}).$
Here, the first equality comes from our definition of $\approx$ from $\approx_{0}$, and the fact that each pair in $\approx_{0}$ has elements contained entirely in a universe $\mathsf{dom}(\mathbf{A}_{i})$; the second equality follows from this fact and our definitions of $\overline{a}$ and $\approx_{0}$; the third equality follows from our assumption that each $(\mathbf{A}_{i},\overline{t_{i}},\approx_{i})$ is an extended $p_{i}$-representation; the last equality follows from the definition of $\mathsf{out}$.
By induction, each given extended $p_{i}$-representation has a $p_{i}$-decomposition $E_{i}=(T_{i},\chi_{i})$, where $r_{i}$ is the root node of $T_{i}$, with associated mappings $\alpha_{i},\beta_{i}$. We assume without loss of generality that the vertex sets $V(T_{i})$ are pairwise disjoint (if not, they can be renamed to achieve this). We provide a $p$-decomposition as follows. Combining the trees $T_{i}$, create a rooted tree $T$ with one extra vertex $r$ whose children are $r_{1},\ldots,r_{\ell}$. Define $\alpha:\mathsf{subplans}(p)\to V(T)$ to be the mapping that extends each mapping $\alpha_{i}$ and where $\alpha(p)=r$; define $\beta:\mathsf{subplans}(p)\to\mathsf{dom}(\mathbf{A})\langle\approx\rangle^{*}$ to be the mapping with $\beta(p)=\overline{a}\langle\approx\rangle$, and where, for each subplan $q$ of a plan $p_{i}$, when $\beta_{i}(q)=(c_{1},\ldots,c_{m_{i}})$, $\beta(q)$ is defined as the tuple obtained by further dividing each entry by $\approx^{*}$, that is, as the tuple $(c^{\prime}_{1},\ldots,c^{\prime}_{m_{i}})$ where each $c^{\prime}_{j}$ is the unique $\approx^{*}$-equivalence class containing $c_{j}$; such a class exists, since each $c_{j}$ is a $\approx_{i}^{*}$-equivalence class, and $\approx_{i}^{*}$ is contained in $\approx^{*}$ (due to $\approx_{i}$ being contained in $\approx$).
Define $\chi(r)=\{\overline{a}\langle\approx\rangle\}$. For each subplan $q$ of $p$ with $q\neq p$, define $\chi({\alpha(q)})=\{\beta(q)\}$. It is straightforward to verify that $(T,\chi)$ is a rooted tree decomposition as required by the definition of $p$-decomposition with mappings $\alpha$ and $\beta$ as defined above.
It remains to show that the $p$-decomposition $(T,\chi)$ satisfies the containment property. We showed that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation, so it holds in the case that $q=p$. Suppose that $q$ is a subplan of $p$ with $q\neq p$. Then $q$ is a subplan of a plan $p_{j}$, for some $j\in[\ell]$. Since $(\mathbf{A}_{j},\overline{t_{j}},\approx_{j})$ is an extended $p_{j}$-representation, we have for any structure $\mathbf{D}$ over $\sigma$ that
$\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},\beta_{j}(q),\mathbf{D}).$
Let us make some observations. First, we have $\alpha(q)=\alpha_{i}(q)$, and so $\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$ is a substructure of $\mathbf{A}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$. We can write $\beta_{j}(q)$ as a tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle$ where $a^{\prime}_{1},\ldots,a^{\prime}_{k}\in\mathsf{dom}(\mathbf{A}_{j})$. By our definition of $\beta(q)$ from $\beta_{j}(q)$, we have $\beta(q)=(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$. We have
$\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},\beta_{j}(q),\mathbf{D})$
$=\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle,\mathbf{D})$
$\supseteq\mathsf{homs}(\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E},(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle,\mathbf{D}).$
The equality follows from the above observations on $\beta_{j}(q)$. The containment can be explained as follows. We have that $\approx_{j}$ is a subset of $\approx$, and, on nodes below $\alpha(q)$, the tree decomposition $E$ is obtained from the tree decomposition $E_{j}$ by further dividing each element in a set associated via $\chi_{j}$ by $\approx^{*}$. We thus have a canonical homomorphism from $\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$ to $\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E}$, namely, the mapping that further divides each element of the first structure by $\approx^{*}$. Since this homomorphism maps the tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle$ to the tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$, the containment holds by Proposition 2.2.
By chaining together the established relationships and using $\beta(q)=(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$, we conclude $\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E},\beta(q),\mathbf{D})$, as required.
We finish this section by showing how the $p$-representation of a plan $p$ can be used to determine the output degree of the plan. To this end, we present two lemmas providing upper and lower bounds on the size of the result of evaluating an open structure on a structure. This lemma, providing an upper bound, is a reformulation of Theorem 4.4 of (Gottlob et al., 2012).
Lemma 4.5.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $\mathbf{D}$ be a structure over $\sigma$ that satisfies $\Sigma$. Let $(\mathbf{A},\overline{a})$ be an open structure over $\sigma$. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$. It holds that
$|\mathsf{homs}(\mathbf{A},\{\overline{a}\},\mathbf{D})|\leq M_{\mathbf{D}}^{C_{\Sigma}^{\mathbf{B}}(\{\overline{b}\})}.$
We next present a complementary lemma that provides the matching lower bound and which again follows directly from Theorem 4.4 of (Gottlob et al., 2012).
Lemma 4.6.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $\mathbf{A}$ be a structure over $\sigma$. For any open structure $(\mathbf{A},\overline{a})$, there exists a constant $c>0$ such that, with $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$, there exists a sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures over $\sigma$ that satisfy $\Sigma$, with $M_{\mathbf{D}_{n}}$ unbounded, and where, for each $n~{}\geq~{}1$,
$|\mathsf{homs}(\mathbf{A},\{\overline{a}\},\mathbf{D}_{n})|\geq cM_{\mathbf{D}_{n}}^{C_{\Sigma}^{\mathbf{B}}(\{\overline{b}\})}.$
Now, Theorem 4.4 guarantees the existence of a $p$-representation for any SPJ plan $p$. Combining the existence of a $p$-representation with the upper and lower bounds given by Lemmas 4.5 and 4.6, respectively, we yield the following proposition pin-pointing the output degree of a plan $p$.
Proposition 4.7.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ be a $p$-representation. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$. Then, $p$ has output degree $C^{\mathbf{B}}_{\Sigma}(\{\overline{b}\})$.
5. Intermediate Size Bounds
We devote this section to formalizing and establishing the statements used to prove Theorem 3.3. We start with statement (i), i.e. we show that we can synthesize a plan which is $\Sigma$-equivalent to the one given originally and which has the desired intermediate degree.
Theorem 5.1 (Plan synthesis).
There exists a polynomial-time algorithm that, given a set $\Sigma$ of unary keys over a signature $\sigma$ and an open structure $(\mathbf{C},\overline{c})$ over $\sigma$ with $\textrm{Chase}_{\Sigma}(\mathbf{C},\overline{c})=(\mathbf{C},\overline{c})$ and which is a $p$-representation for some plan $p$, outputs a well-behaved plan $p^{\prime}$ where (a) $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$, and (b) $p^{\prime}$ has intermediate degree $\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$.
Proof.
The heart of the proof can be derived from argumentation in Section 4 of (Gottlob et al., 2012). As we draw on this section, we employ the terminology of (Gottlob et al., 2012), and refer to their argumentation. In particular, their result refers to conjunctive queries and functional dependencies. The former are known to be equivalent to SPJ queries (and are in syntax close to our $p$-representations), while the latter generalize key constraints. In the following arguments, as in (Gottlob et al., 2012), for elements $X_{i},X_{j}\in\mathsf{dom}(\mathbf{C})$, we use the notation $X_{i}\rightarrow X_{j}$ to say that element $X_{j}$ is functionally determined by $X_{i}$ via some functional dependency. In Section 4 of (Gottlob et al., 2012), it is explained how to transform a conjunctive query $Q_{c}$, to which the chase has been applied, to a conjunctive query $Q^{\prime}$ over a modified set of relations that has no functional dependencies. Let us for the moment put aside the output relation and output variables, and consider the action of this transformation on the input relations and variables. The transformation considers functional dependencies of the form $X_{i}\rightarrow X_{j}$ over pairs of input variables, and iteratively removes them; a removal is performed by adding entailed functional dependencies, and by expanding each atom including $X_{i}$ to a new atom where $X_{j}$ is added (if not already present).
For each functional dependency $X_{i}\rightarrow X_{j}$, we can maintain a binary relation $S_{ij}$ whose first coordinate is a key for the relation, and such that adding $S_{ij}(X_{i},X_{j})$ to the query atoms would not change the query answers; this holds initially by taking $S_{ij}$ to be an appropriate projection of the relation of an atom mentioning both $X_{i}$ and $X_{j}$, and then when an entailed dependency $X_{k}\rightarrow X_{j}$ is added, due to dependencies $X_{k}\rightarrow X_{i}$ and $X_{i}\rightarrow X_{j}$, we can derive $S_{kj}$ (if needed) by taking the natural join of $S_{ki}$ and $S_{ij}$ and then projecting the middle coordinate. It is readily verified that the sizes of these relations $S_{ij}$ never exceeds the size of any input relation.
When an atom including a variable $X_{i}$ is expanded to include the variable $X_{j}$, a new relation for the atom can be computed by joining the original relation with the relation $S_{ij}$. Since each relation $S_{ij}$ expresses a functional dependency, this never increases the size of any relation.
The result of the transformation, applied to the input atoms of a conjunctive query, is thus a new set of input atoms where every new relation can be derived from the original relations via a well-behaved SPJ plan of intermediate degree $1$, where every new relation has size bounded above by an original relation, and where the answers (over all variables) to the new set of atoms is the same as the answers to the original set. Consider the application of the transformation to the input atoms of the conjunctive query corresponding to $\mathbf{C}$. Let $\mathbf{C}^{\prime}$ be the relational structure representing the new set of atoms. It is moreover shown in Lemma 4.7 of (Gottlob et al., 2012) that the color number of an overall query (including input variables) is not changed by this transformation, and this implies that $C^{\mathbf{C}^{\prime}}_{\Sigma}(S)\leq C^{\mathbf{C}^{\prime}}_{\Sigma}(S^{\prime})=C^{\mathbf{C}}_{\Sigma}(S)$ for any set $S$ of variables and its expansion $S^{\prime}$ after applying the transformation. Let us then consider $(\mathbf{C}^{\prime},\overline{c})$; due to the just-given inequality, we have $C_{\Sigma}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$, and it thus suffices to establish the result for the empty set of unary keys $\Sigma^{\prime}=\emptyset$ along with $(\mathbf{C}^{\prime},\overline{c})$, in place of $\Sigma$ with $(\mathbf{C},\overline{c})$.
To establish the result for $(\mathbf{C}^{\prime},\overline{c})$ in the absence of keys, it suffices to compute a tree decomposition $F$ of $(\mathbf{C}^{\prime},\overline{c})$ where $C_{\Sigma}^{\mathbf{C}^{\prime}}\textup{-width}(F)$ is equal to $C_{\Sigma}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})$, which can be done in exponential time, and then compute a plan that materializes the answers to each bag, where the answers for a bag with variables $S$ would be considered to be the join of the set of atoms obtained from $\mathbf{C}^{\prime}$’s conjunctive query by taking each atom, and projecting onto $S$. These different sets of answers can then be combined via joins and projects in a standard way, following the structure of the tree decomposition. But to find a plan that does this materialization for a bag with variables $\{a_{1},\ldots,a_{n}\}$, we simply follow the proof of Theorem 6 of (Atserias et al., 2013), which explains how to construct a join-project plan for any pure join plan; in doing so, they define plans $\phi_{1},\ldots,\phi_{n}$; each of these plans is implementable as a well-behaved one, since they iteratively introduce one variable at a time: in their notation, the plan $\phi_{i}$, which can be viewed as a multiway join, concerns variables $\{a_{1},\ldots,a_{i}\}$.
We continue to show that statement (ii) of Section 3 also holds. To that end, let us first formalize the statement. In essence, it provides a lower bound on the intermediate degree of any plan that is $\Sigma$-semantically equivalent to $p$.
Theorem 5.2.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ be a $p$-representation. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$, and let $(\mathbf{C},\overline{c})$ be a core of $(\mathbf{B},\overline{b})$. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. For any SPJ plan $p^{+}$ that is $\Sigma$-semantically equivalent to $p$, it holds that $p^{+}$ has a subplan with output degree $\geq d$.
Before proceeding with the proof, we establish two preparatory lemmas that will aid us in showing the lower bound on the intermediate degree. First, we show that for an open structure $(\mathbf{C},\overline{c})$ which is a core of the chase of a $p$-representation and which has width $d$, there is a vertex in any tree decomposition of $(\mathbf{C},\overline{c})$ witnessing that the intermediate results grow in accordance with a degree $d$ polynomial.
Lemma 5.3.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ be a $p$-representation. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$, and let $(\mathbf{C},\overline{c})$ be a core of $(\mathbf{B},\overline{b})$. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. For any tree decomposition $(T,\chi)$ of $(\mathbf{C},\overline{c})$, there exists a vertex $t_{0}$ of $T$, a constant $c^{\prime}>0$, and a sequence of structures $(\mathbf{D}_{n})_{n\geq 1}$ with $M_{\mathbf{D}_{n}}$ unbounded where $|\mathsf{homs}(\mathbf{C},\chi(t_{0}),\mathbf{D}_{n})|\geq c^{\prime}M_{\mathbf{D}_{n}}^{d}$.
Proof.
[…]
We next give a lemma whose assumptions are a relaxation of (that is, they are implied by) the condition that $\mathbf{C}$ is a core of $\mathbf{B}$. It allows us to lower bound, for any subset $S\subseteq\mathsf{dom}(\mathbf{B})$, the size of $\mathsf{homs}(\mathbf{B},S,\cdot)$ by the size of $\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\cdot)$.
Lemma 5.4.
Suppose that $\mathbf{B},\mathbf{C}$ are structures over signature $\sigma$ such that $\mathbf{C}$ is a substructure of $\mathbf{B}$, and there exists a retraction $h$ from $\mathbf{B}$ to $\mathbf{C}$. Then, for any subset $S\subseteq\mathsf{dom}(\mathbf{B})$ and any structure $\mathbf{D}$ over $\sigma$, it holds that $|\mathsf{homs}(\mathbf{B},S,\mathbf{D})|\geq|\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\mathbf{D})|$.
Proof.
We show how to pass from any map in $\mathsf{homs}(\mathbf{C},S~{}\cap~{}\mathsf{dom}(\mathbf{C}),\mathbf{D})$ to a map in $\mathsf{homs}(\mathbf{B},S,\mathbf{D})$, and then explain why the passage is injective, which suffices. Suppose that $g\in\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\mathbf{D})$ and $h$ is a retraction from $\mathbf{B}$ to $\mathbf{C}$. Then there exists an extension $g^{\prime}$ of $g$ that is in $\mathsf{homs}(\mathbf{C},\mathsf{dom}(\mathbf{C}),\mathbf{D})$, from which it follows that the composition $g^{\prime}\circ h$ is in $\mathsf{homs}(\mathbf{B},\mathsf{dom}(\mathbf{B}),\mathbf{D})$ (in writing this composition, $h$ is applied first). It follows that the restriction $(g^{\prime}\circ h)\upharpoonright S$ is in $\mathsf{homs}(\mathbf{B},S,\mathbf{D})$. This passage is injective since if two mappings in $\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\mathbf{D})$ differ on an element $c$ of $S\cap\mathsf{dom}(\mathbf{C})$, the two mappings that they are passed to will also differ on $c$, due to the assumption that $h$ acts as the identity on $C$.
We proceed to prove our final theorem.
Proof of Theorem 5.2.
Fix a set $\Sigma$ of unary keys and a SPJ plan $p$ over $\Sigma$. Let $(\mathbf{A},\overline{a})$ be a $p$-representation, $(\mathbf{B},\overline{b})=\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, and $(\mathbf{C},\overline{c})$ be a core of $(\mathbf{B},\overline{b})$. For any structure $\mathbf{D}$ satisfying $\Sigma$ we have that
$\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})=\mathsf{homs}(\mathbf{B},\overline{b},\mathbf{D})=\mathsf{homs}(\mathbf{C},\overline{c},\mathbf{D}),$
where the first equality holds by definition of $p$-representation, the second equality holds by definition of the chase procedure and the third equality holds by Propositions 2.4.
Let $p^{+}$ be any SPJ plan that is $\Sigma$-semantically equivalent to $p$. Let $(\mathbf{A}^{+},\overline{a^{+}})$ be a $p^{+}$-representation and $(\mathbf{B}^{+},\overline{b^{+}})=\textrm{Chase}_{\Sigma}(\mathbf{A}^{+},\overline{a^{+}})$. For any structure $\mathbf{D}$ satisfying $\Sigma$ we have that
$\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{+},\mathbf{D})=\mathsf{homs}(\mathbf{A}^{+},\overline{a^{+}},\mathbf{D})=\mathsf{homs}(\mathbf{B}^{+},\overline{b^{+}},\mathbf{D}),$
where the first equality holds by $\Sigma$-semantic equivalence and the second and third equality hold by definition as above.
Since we have that $\mathsf{homs}(\mathbf{C},\overline{c},\mathbf{D})=\mathsf{homs}(\mathbf{B}^{+},\overline{b^{+}},\mathbf{D})$ for all structures $\mathbf{D}$ satisfying $\Sigma$, and we know that both $\mathbf{C}$ and $\mathbf{B}^{+}$ are structures satisfying $\Sigma$, it holds specifically that $\overline{c}\in\mathsf{homs}(\mathbf{C},\overline{c},\mathbf{C})=\mathsf{homs}(\mathbf{B}^{+},\overline{b^{+}},\mathbf{C})$ and $\overline{b^{+}}\in\mathsf{homs}(\mathbf{B}^{+},\overline{b^{+}},\mathbf{B}^{+})=\mathsf{homs}(\mathbf{C},\overline{c},\mathbf{B}^{+})$. This immediately implies that there is a homomorphism from $(\mathbf{B}^{+},\overline{b^{+}})$ to $(\mathbf{C},\overline{c})$, and vice-versa, showing that the two open structures are homomorphically equivalent. Since $(\mathbf{C},\overline{c})$ is a core, by Proposition 2.4 we have that $\mathsf{aug}(\mathbf{C},\overline{c})$ is isomorphic to a substructure of $\mathsf{aug}(\mathbf{B}^{+},\overline{b^{+}})$. By renaming the elements of $\mathsf{dom}(\mathbf{B}^{+})$, we can achieve that $\mathsf{aug}(\mathbf{C},\overline{c})$ is a core of (and thus a substructure of) $\mathsf{aug}(\mathbf{B}^{+},\overline{b^{+}})$. We then have that $\overline{b^{+}}=\overline{c}$.
By Theorem 4.4, there exists a $p^{+}$-decomposition $E^{+}=(T,\chi^{+})$ of $(\mathbf{A}^{+},\overline{a^{+}})$ with associated mappings $\alpha,\beta$. Within $E^{+}$, we equate elements following the chase procedure on $(\mathbf{A},\overline{a})$ such that the resulting $E^{+}$ is a decomposition of $(\mathbf{B}^{+},\overline{b^{+}})$. We define a tree decomposition of $(\mathbf{C},\overline{c})$ as follows. Take the pair $(T,\chi)$ where $T$ is as in $E^{+}$, and where, for each vertex $t$ of $T$, we define $\chi(t)=\chi^{+}(t)\cap\mathsf{dom}(\mathbf{C})$. It is straightforward to verify that $(T,\chi)$ is a tree decomposition of $(\mathbf{C},\overline{c})$. By Lemma 5.3, there exists a vertex $t$ of $T$, a constant $c^{\prime}>0$, and structures $(\mathbf{D}_{n})_{n\geq 1}$ with $M_{\mathbf{D}_{n}}$ unbounded, where $|\mathsf{homs}(\mathbf{C},\chi(t),\mathbf{D}_{n})|\geq c^{\prime}M^{d}_{\mathbf{D}_{n}}$. There exists a subplan $q^{+}$ of $p^{+}$ such that $\alpha(q^{+})=t$. We then have
$|\mathsf{out}(q^{+},\mathbf{D}_{n})|$ $\geq|\mathsf{homs}(\mathbf{B}_{\leq t,E^{+}},\chi^{+}(t),\mathbf{D}_{n})|$
$\geq|\mathsf{homs}(\mathbf{B},\chi^{+}(t),\mathbf{D}_{n})|$
$\geq|\mathsf{homs}(\mathbf{C},\chi(t),\mathbf{D}_{n})|$
$\geq c^{\prime}M^{d}_{\mathbf{D}_{n}}.$
The first inequality holds by the definition of $p^{+}$-decomposition; the second inequality holds since $\mathbf{B}_{\leq t,E^{+}}$ is a substructure of $\mathbf{B}$; the third inequality follows from Lemma 5.4, whose assumptions hold because $\mathsf{aug}(\mathbf{C},\overline{c})$ is a core of $\mathsf{aug}(\mathbf{B},\overline{b})$; the last inequality follows from our choice of $(\mathbf{D}_{n})_{n\geq 1}$. It follows that the subplan $q^{+}$ has output degree $\geq d$.
6. Conclusion
The intermediate degree asymptotically characterizes the size of intermediate relations obtained during the execution of a query plan. For SPJ plans, we have shown how to optimize for this measure among all semantically equivalent SPJ plans with respect to a given set of unary key constraints. A natural way to extend our work is to consider either more expressive query languages, by adding more operators, or more expressive constraints, such as arbitrary functional dependencies.
Appendix A Proof of Proposition 2.2
For the forward direction, let $h$ be a homomorphism as described; then, for any homomorphism $g$ from $\mathbf{A}^{\prime}$ to $\mathbf{D}$, the composition $g\circ h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$ where $g(\overline{a^{\prime}})=g(h(\overline{a}))$. For the backward direction, set $\mathbf{D}=\mathbf{A}^{\prime}$; then, we have that $\overline{a^{\prime}}\in\mathsf{homs}(\mathbf{A}^{\prime},\overline{a^{\prime}},\mathbf{D})$, implying that $\overline{a^{\prime}}\in\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$, and thus there exists a homomorphism from $\mathbf{A}$ to $\mathbf{A}^{\prime}$ sending $\overline{a}$ to $\overline{a^{\prime}}$.
Appendix B Proof of Lemma 3.5
Suppose that there existed a value $\epsilon>0$ such that a plan $p$ had output degree $\geq d$ and output degree $\leq(d-\epsilon)$. Then, there would be functions $g\in\Omega(M^{d})$ and $h\in O(M^{d-\epsilon})$ where $g(M_{\mathbf{D}_{n}})\leq|\mathsf{out}(p,\mathbf{D}_{n})|\leq h(M_{\mathbf{D}_{n}})$ for all structures $\mathbf{D}_{n}$ in a sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures with $M_{\mathbf{D}_{n}}$ unbounded, implying that $g\leq h$ on infinitely many values, a contradiction.
Appendix C Proof of Lemma 3.6
The forward direction is immediate from the definitions. For the backward direction, we have that, for each subplan $q$ of $p$, there exists a function $h_{q}(M)\in O(M^{d})$ where, for all structures $\mathbf{D}$ over the signature of $p$, $|\mathsf{out}(q,\mathbf{D})|\leq h_{q}(M_{\mathbf{D}})$. The plan $p$ then has intermediate degree $\leq d$ via the function $\sum_{q}h_{q}$, where the sum is over all subplans $q$ of $p$.
Appendix D Proof of Omitted Cases of Theorem 4.4
Suppose that $p=R$. Let $m$ be the arity of $R$. Define $\mathbf{A}$ as the structure with $A=\{a_{1},\ldots,a_{m}\}$, $R^{\mathbf{A}}=\{(a_{1},\ldots,a_{m})\}$, and $S^{\mathbf{A}}=\emptyset$ for each $S\in\sigma\setminus\{R\}$. Assign $\approx$ to be the empty set; we have that $\approx^{*}$ is the equality relation on $\mathsf{dom}(\mathbf{A})$. Letting $\overline{a}$ denote $(a_{1},\ldots,a_{m})$, we have that $(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle)$ is isomorphic to $(\mathbf{A},\overline{a})$. It is evident that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation.
We give a $p$-decomposition of this representation as follows. Let $T$ be the tree with one vertex $t$, and define $\chi(t)=\{\overline{a}\}$. We then have that $(T,\chi)$ is a $p$-decomposition, via the mapping $\alpha$ with $\alpha(p)=t$ and the mapping $\beta$ with $\beta(p)=\overline{a}$. The containment property is, in this case, a direct consequence of the established fact that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation.
Suppose that $p$ has the form $\pi_{j_{1},\ldots,j_{n}}(p_{0})$, where $p_{0}$ has arity $m$ and $j_{1},\ldots,j_{n}\in[m]$. By induction there exist an extended $p_{0}$-representation $(\mathbf{A},(a_{1},\ldots,a_{m}),\approx)$ and a $p_{0}$-decomposition $(T_{0},\chi_{0})$ thereof with root $r_{0}$ and mappings $\alpha_{0}$, $\beta_{0}$. Define $\overline{a}$ to be the tuple $(a_{j_{1}},\ldots,a_{j_{n}})$; we then have that for any structure $\mathbf{D}$ over $\sigma$:
$\mathsf{homs}(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle,\mathbf{D})=\pi_{j_{1},\ldots,j_{n}}(\mathsf{homs}(\mathbf{A}\langle\approx\rangle,(a_{1},\ldots,a_{m})\langle\approx\rangle,\mathbf{D}))$
and thus that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation.
We obtain a $p$-decomposition $(T,\chi)$ of this representation by extending $(T_{0},\chi_{0})$ as follows: create a root node $r$ whose unique child is $r_{0}$, and let $\chi(t)=\chi_{0}(t)$ for every $t\in T_{0}$ and define $\chi(r)=\{\overline{a}\langle\approx\rangle\}$. Extend the mappings $\alpha_{0}$, $\beta_{0}$ to mappings $\alpha,\beta$ by defining $\alpha(p)=r$ and $\beta(p)=\overline{a}\langle\approx\rangle$. The containment property on each subplan of $p_{0}$ holds as a direct consequence of it having held on the $p_{0}$-decomposition. When we consider $p$ (as a subplan of $p$), the containment property holds due to $(\mathbf{A},\overline{a},\approx)$ being an extended $p$-representation.
Appendix E Proof of Proposition 4.7
The existence of a $p$-representation follows from Theorem 4.4. Fix some structure $\mathbf{D}$ over $\sigma$. As $(\mathbf{A},\overline{a})$ is a $p$-representation, we have that $\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$. It follows from Lemma 4.5 that $|\mathsf{out}(p,\mathbf{D})|\leq M_{\mathbf{D}}^{d}$, and so $p$ has output degree $\leq d$. By Lemma 4.6, there is a constant $c>0$ and an infinite sequence of structures $(\mathbf{D}_{n})_{n\geq 1}$, with $M_{\mathbf{D}_{n}}$ unbounded, where $|\mathsf{out}(p,\mathbf{D}_{n})|\geq cM_{\mathbf{D}_{n}}^{d}$. We obtain that the plan $p$ has output degree $\geq d$, and can thus conclude by Lemma 3.5 that $p$ has output degree $d$.