### v2

# Intermediate Relation Size Bounds for Select-Project-Join-Union Query Plans

Abstract
We study the problem of statically optimizing select-project-join-union (SPJU) plans where unary key constraints are allowed. A natural measure of a plan, which we call the output degree and which has been studied previously, is the minimum degree of a polynomial bounding the plan’s output relation, as a function of the input database’s maximum relation size. This measure is, by definition, invariant under passing from a plan to another plan that is semantically equivalent to the first. In this article, we consider a plan measure which we call the intermediate degree; this measure is defined to be the minimum degree of a polynomial bounding the size of all intermediate relations computed during a plan’s execution — again, as a function of the input database’s maximum relation size. We present an algorithm that, given an SPJU plan $p$ and a set $\Sigma$ of unary keys, computes an SPJU plan $p^{\prime}$ that is semantically equivalent to $p$ (over databases satisfying $\Sigma$) and that has the minimum intermediate degree over all such semantically equivalent plans. For the types of plans considered, we thus obtain a complete and effective understanding of intermediate degree.
keywords:
Select-Project-Join-Union Plans, Query Optimization, Tree Decompositions, Key Dependencies
1 Introduction
Background and motivation. Modern database systems capable of evaluating queries of the standard query language SQL typically have, as a key component, a query optimizer that synthesizes execution plans consisting of basic operations from a relational algebra; techniques for estimating and comparing the costs of evaluating such plans are considered central in the study of databases [1, Chapter 6]. In this article, we study a basic measure of execution plans: the intermediate size, by which we mean the maximum size over all intermediate relations computed during a query’s execution, relative to a database. We view this as a natural and fundamental measure: the intermediate size gives an obvious lower bound on the amount of time needed to execute a plan, for algorithms that explicitly materialize all intermediate relations; and, it characterizes the amount of space needed to store the results of all such intermediate relations.
We study this size measure in an asymptotic setting, as a function of the maximum relation size $M_{\mathbf{D}}$ of a database $\mathbf{D}$. We say that a plan has intermediate degree $d$ when $d\geq 0$ and there exists a degree $d$ polynomial upper bounding the plan’s intermediate size, as a function of $M_{\mathbf{D}}$, but no lower degree polynomial provides such a bound. We study this measure for SPJU query plans, which allow select, project, join and union as operations on relations. These plans are semantically equivalent to unions of conjunctive queries, and are intensely studied throughout database theory. We permit unary key constraints to be expressed on database relations. In essence, a unary key constraint states that there is a single coordinate of a relation where, for each tuple in the relation, knowing the value at the coordinate determines the entire tuple.
Having identified a relevant complexity measure of query plans—the intermediate degree—and a relevant class of query plans—the SPJU query plans, one of the most basic and natural questions one could ask, from the standpoint of query optimization, is this: is there an algorithm that receives a query plan as input, and reformulates it into a semantically equivalent query plan whose intermediate degree is minimal, over all such equivalent plans? That is, is there an algorithm that can statically analyze a given query plan, and reformulate it so as to optimally minimize the intermediate degree, without disrupting the plan’s semantics? A primary contribution of this article is a positive answer to this question.
Contributions. Our main theorem presents a minimization algorithm that, given an SPJU plan $p$ with a set $\Sigma$ of unary key constraints, computes an SPJU plan $p^{\prime}$ that is semantically equivalent to $p$ (with respect to databases satisfying the constraints $\Sigma$), and has the lowest possible intermediate degree over all SPJU plans that are so semantically equivalent to $p$. This theorem thus yields, in the asymptotic setting considered, a full understanding of how to statically optimize the intermediate degree so that it is the minimum possible. We remark that (to our knowledge) this type of understanding was not previously known even for SPJ plans (SPJU plans that do not use the union operator) and on databases without any constraints. We view the definition of the notion of intermediate degree, and its supporting theory, as conceptual contributions of this work.
Our result showing how to minimize the intermediate degree amounts to demonstrating how to compute, for any given plan, an asymptotically optimal procedure in a restricted model of computation, namely, the class of SPJU plans. Our focus on polynomial degree is aligned with the notion of fixed-parameter tractability, whereby a problem is tractable essentially when there is a degree $d$ such that each so-called slice has an algorithm with running time bounded above by a degree $d$ polynomial; a slice is a set of all instances sharing the same parameter, which is a value associated with each instance.
To prove our result, we make a number of technical contributions that we believe will be of significant utility in the future study of intermediate degree, which we view as a highly natural measure of plan complexity. One tool that we present (in Subsection 4.1) is a theorem showing how to convert an SPJ query plan to a relational structure along with a certain form of tree decomposition of the structure; this allows us to apply structural measures based on tree decompositions to reason about the intermediate degree of plans. As a first step towards our main result, we show a restricted version of the main theorem to SPJ plans, that is, an algorithm that, given an SPJ plan $p$ and a set $\Sigma$ of unary keys, outputs a semantically equivalent SPJ plan $p^{\prime}$ (over all databases satisfying $\Sigma$) with the lowest possible intermediate degree. While the way that this algorithm computes the minimized plan $p^{\prime}$ from a given plan $p$ is, on a high level, quite natural — essentially, $p^{\prime}$ is based on the core of the chase of a relational structure naturally derived from $p$ — the proof that $p^{\prime}$ is minimal in the desired sense is non-trivial, since we have to reason about arbitrary plans that are semantically equivalent to $p^{\prime}$. Intuitively, the core is a minimized version of a relational structure and the chase is a well-known procedure that enforces constraints on a relational structure.
As defined, the intermediate degree is a measure of query plans. It follows from our definitions that any algorithm for evaluating a query plan $q$ that materializes all intermediate relations must require time at least $\Omega(M_{\mathbf{D}}^{d})$, where $d$ is the intermediate degree of $q$. That is, the polynomial time dependence on $M_{\mathbf{D}}$ must be of degree at least $d$. Having observed this lower bound, we are naturally led to the question of whether or not this bound is tight. We present results indicating that it is, which renders the intermediate degree as an algorithmically meaningful measure. When given a plan $p$, our minimization algorithm can always output a plan $p^{\prime}$ with the properties described above, and with the additional syntactic property that $p^{\prime}$ is well-behaved; as we show, well-behavedness of a plan $q$ implies that there is an evaluation algorithm for evaluating $q$ on a database $\mathbf{D}$ that always runs within time $M_{\mathbf{D}}^{d}$ times a multiplicative overhead, where $d$ is the intermediate degree of $q$. So, a well-behaved plan can be algorithmically evaluated in time close to $O(M_{\mathbf{D}}^{d})$, complementing the time lower bound of $\Omega(M_{\mathbf{D}}^{d})$ discussed above. We leave an analysis of the exact nature of the multiplicative overhead as a question for future work; we wish to emphasize here that the analysis of our evaluation algorithm and the resulting time bound are obtained in an elementary fashion.
Related Work. Our work is inspired by and is closely related to the work of Atserias, Grohe, and Marx [12, 2], as well as the follow-up work thereof by Gottlob, Lee, Valiant, and Valiant [10]. A main result shown by [2] yields a general upper bound—the AGM bound—on the output size of any pure join plan $q$, namely, a bound of $M_{\mathbf{D}}^{\rho^{*}(q)}$, where $\rho^{*}(q)$ is the so-called fractional edge cover number of the plan $q$. They also supply an argument that this upper bound is tight. Here, by a pure join plan, we mean a plan that consists only of joins and basic relations of a database.
Let us define the output degree of a query to be the minimum degree of a polynomial bounding the query’s output relation, as a function of the maximum relation size $M_{\mathbf{D}}$ of a database $\mathbf{D}$. Note that while the output degree is clearly invariant under passing from a query to another query that is semantically equivalent to the first, the intermediate degree does not enjoy this invariance. The AGM bound along with the argument that it is tight establish $\rho^{*}(q)$ as the output degree of a pure join plan $q$. The article [2] also exhibits, for each such pure join plan, a semantically equivalent join-project plan whose intermediate relations all have size less than (or equal to) the output size, implying a characterization of the minimum intermediate degree (over semantically equivalent plans) as the fractional edge cover number—for pure join plans. We view the article [2] as begging the study of more general classes of plans: they show explicitly the necessity of using projections in the plans witnessing that the intermediate relation sizes can be made to be at most the output size!
Extending the AGM bound, the work [10] characterizes the output degree of any SPJ plan and considering databases where unary key constraints can be expressed; this is done by identifying a measure called the color number. Note that this measure generalizes the fractional edge cover number (as discussed in Section 4 of [10]). Our work builds on this measure and the accompanying understanding of unary key constraints. In particular, in the present work, we naturally extend the notion of color number to define a measure of tree decompositions. We remark that the results of [10] are phrased in terms of conjunctive queries, which are semantically equivalent to SPJ plans.
An important and influential line of research emerging after the work of Atserias, Grohe, and Marx presented new join algorithms, so-called worst-case optimal join algorithms, with running times shown to be bounded above by the AGM bound—apart from a multiplicative overhead (see for example [19, 18, 17] and the discussion therein). Building on this work, further research presented algorithms, with running time analyses, for more general forms of queries, including queries with aggregates [13, 14]. The focus of these further works are on presenting algorithms and upper bounding their running time in terms of structural measures, generally without considering how to minimize queries up to semantic equivalence; the present article establishes a complementary result by showing, for each query plan $p$, a tight lower bound on the running time of a restricted class of procedures, namely, procedures that can be described by plans that are semantically equivalent to the given plan $p$.
Perspective. One measure of query plans that has been studied heavily is width, which by now has a mature underlying theory (refer for example to [15, 11, 6, 3, 20, 4, 7]); while this measure is often defined on logical formulas (and is sometimes formulated as the number of variables), for a query plan the width amounts to the maximum arity over all subplans. The present article is a contribution to a general research direction whereby we hope to identify further meaningful and informative measures of query plans, and study how to minimize various classes of queries with respect to introduced measures.
Outline. In Section 2, we present the necessary concepts used throughout the paper. Section 3 defines the central notion of intermediate degree, gives an intuition for the main theorem as well as discusses some algorithmic aspects of the plan output by the algorithm underlying the main theorem. In Section 4 we show how we can prove the main theorem restricted to SPJ plans before generalizing our result to SPJU plans in Section 5.
2 Preliminaries
We use $\mathbb{Q}$ to denote the rational numbers, $\mathbb{Q}^{+}$ to denote the non-negative rational numbers, and $\mathbb{N}$ to denote the natural numbers, which we understand to include 0. For a natural number $k\in\mathbb{N}$, we use $[k]$ to denote the first $k$ positive natural numbers, that is, the set $\{i\in\mathbb{N}\penalty 10000\ |\penalty 10000\ 1\leq i\leq k\}$. For a map $f:A\to B$ and some $S\subseteq A$, $f\upharpoonright S$ denotes the restriction of $f$ to $S$. For a tuple $\overline{a}=(a_{1},\ldots,a_{k})$ over a set $A$ and a map $f:A\to B$, by extension of notation, we write $f(\overline{a})$ for the entry-wise action of $f$ on the tuple, i.e., to denote $(f(a_{1}),\ldots,f(a_{k}))$; we further write $\{\overline{a}\}$ for the set $\{a_{1},\ldots,a_{k}\}$. When $A$ is a set, we use $A^{*}$ to denote the set of all finite-length tuples over $A$. When $\approx$ is a binary relation on a set $A$, we define $\approx^{*}$ as the reflexive-symmetric-transitive closure of $\approx$, that is, $\approx^{*}$ is the smallest equivalence relation on $A$ that contains $\approx$ as a subset.
2.1 Structures
A signature $\sigma$ is a finite set of relation symbols, each with an associated arity. We denote by $R/n$ that symbol $R$ has arity $n\in\mathbb{N}$; we may also write $\mathrm{ar}(R)$ for $n$. A structure $\mathbf{A}$ over a signature $\sigma$ consists of a finite set called the universe or domain of the structure, denoted $\mathsf{dom}(\mathbf{A})$, and, for each symbol $R\in\sigma$, a relation $R^{\mathbf{A}}$ over $\mathsf{dom}(\mathbf{A})$ whose arity is that of $R$. An isolated element of a structure $\mathbf{A}$ is a universe element $a\in\mathsf{dom}(\mathbf{A})$ such that $a$ does not appear in any tuple of any relation of $\mathbf{A}$. An open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$ is a pair consisting of a structure $\mathbf{A}$ over signature $\sigma$ with no isolated elements and a tuple $\overline{a}$ over the universe $\mathsf{dom}(\mathbf{A})$ of $\mathbf{A}$. For a structure $\mathbf{A}$ over signature $\sigma$, let $M_{\mathbf{A}}$ be the maximum size over all relations of $\mathbf{A}$, that is, $\max_{R\in\sigma}|R^{\mathbf{A}}|$.
Two (open) structures are similar if they are defined over the same signature. When $\mathbf{A}$ and $\mathbf{B}$ are similar structures over the signature $\sigma$, we define their union, denoted by $\mathbf{A}\cup\mathbf{B}$, as the structure with universe $\mathsf{dom}(\mathbf{A})\cup\mathsf{dom}(\mathbf{B})$ and where, for each symbol $R\in\sigma$, it holds that $R^{\mathbf{A}\cup\mathbf{B}}=R^{\mathbf{A}}\cup R^{\mathbf{B}}$. For two similar structures $\mathbf{A}$ and $\mathbf{B}$ over a signature $\sigma$, a homomorphism from $\mathbf{A}$ to $\mathbf{B}$ is a map $h:\mathsf{dom}(\mathbf{A})\to\mathsf{dom}(\mathbf{B})$ such that for each symbol $R\in\sigma$, it holds that $\overline{a}\in R^{\mathbf{A}}$ implies $h(\overline{a})\in R^{\mathbf{B}}$. For similar open structures $(\mathbf{A},\overline{a})$ and $(\mathbf{B},\overline{b})$ with $\overline{a}$ and $\overline{b}$ having the same arity, a homomorphism $h$ from $(\mathbf{A},\overline{a})$ to $(\mathbf{B},\overline{b})$ is a homomorphism from $\mathbf{A}$ to $\mathbf{B}$ such that $h(\overline{a})=\overline{b}$. Two similar structures $\mathbf{A}$ and $\mathbf{B}$ are homomorphically equivalent if there exist a homomorphism from $\mathbf{A}$ to $\mathbf{B}$ and a homomorphism from $\mathbf{B}$ to $\mathbf{A}$. An isomorphism from $\mathbf{A}$ to $\mathbf{B}$ is a bijection $h:\mathsf{dom}(\mathbf{A})\to\mathsf{dom}(\mathbf{B})$ such that $h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{B}$, and $h^{-1}$ is a homomorphism from $\mathbf{B}$ to $\mathbf{A}$. Let $(\mathbf{A},\overline{a})$ be an open structure and let $\mathbf{D}$ be a structure that is similar to $\mathbf{A}$; we use $\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$ to denote the relation
$\{h(\overline{a})\penalty 10000\ |\penalty 10000\ \mbox{$h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$}\}.$
We conceive of this relation as the result of evaluating the open structure on the structure $\mathbf{D}$. Similarly, when $\mathbf{A}$ and $\mathbf{D}$ are similar structures and $S\subseteq\mathsf{dom}(\mathbf{A})$, we use $\mathsf{homs}(\mathbf{A},S,\mathbf{D})$ to denote the set of maps
$\{h\upharpoonright S\penalty 10000\ |\penalty 10000\ \mbox{$h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$}\}.$
The next fact follows from the classic work of Chandra and Merlin [5].
Proposition 2.1 ([5]).
Let $(\mathbf{A},\overline{a})$ and $(\mathbf{A}^{\prime},\overline{a^{\prime}})$ be open structures over some signature $\sigma$. There exists a homomorphism from $(\mathbf{A},\overline{a})$ to $(\mathbf{A}^{\prime},\overline{a^{\prime}})$ if and only if for each structure $\mathbf{D}$ over $\sigma$ we have that $\mathsf{homs}(\mathbf{A}^{\prime},\overline{a^{\prime}},\mathbf{D})\subseteq\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$.
For similar structures $\mathbf{A}$ and $\mathbf{B}$, we say that $\mathbf{A}$ is a substructure of $\mathbf{B}$ if $\mathsf{dom}(\mathbf{A})\subseteq\mathsf{dom}(\mathbf{B})$ and for each symbol $R\in\sigma$, it holds that $R^{\mathbf{A}}\subseteq R^{\mathbf{B}}$. When $\mathbf{B}$ is a structure and $S\subseteq\mathsf{dom}(\mathbf{B})$, the induced substructure $\mathbf{B}[S]$ is defined as the structure with universe $S$ and where, for each symbol $R$, it holds that $R^{\mathbf{B}[S]}=\{(b_{1},\ldots,b_{k})\in R^{\mathbf{B}}\penalty 10000\ |\penalty 10000\ \{b_{1},\ldots,b_{k}\}\subseteq S\}$. When $\mathbf{A}$ is a substructure of $\mathbf{B}$, a retraction from $\mathbf{B}$ to $\mathbf{A}$ is a homomorphism from $\mathbf{B}$ to $\mathbf{A}$ that acts as the identity on $\mathsf{dom}(\mathbf{A})$. When there exists a retraction from $\mathbf{B}$ to $\mathbf{A}$, we say that $\mathbf{B}$ retracts to $\mathbf{A}$. A core of a structure $\mathbf{B}$ is a substructure $\mathbf{C}$ of $\mathbf{B}$ such that $\mathbf{B}$ retracts to $\mathbf{C}$, but $\mathbf{C}$ does not retract to any proper substructure of $\mathbf{C}$, that is, any substructure of $\mathbf{C}$ whose universe is a proper subset of $\mathbf{C}$’s universe. A structure is a core if it is a core of itself. We provide an example illustrating the notions of (open) structure, the relation $\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$, substructure, and core for the interested reader in the appendix.
The following facts are known and straightforward to verify.
Proposition 2.2.
The following facts hold.
  1. 1.
    For each structure $\mathbf{B}$, there exists a core of $\mathbf{B}$.
  2. 2.
    A core $\mathbf{C}$ of a structure $\mathbf{B}$ is homomorphically equivalent to $\mathbf{B}$.
  3. 3.
    Any two cores of a structure are isomorphic. (We thus, by a mild abuse of terminology, speak of the core of a structure.)
  4. 4.
    When $\mathbf{C}$ is a core, and $\mathbf{B}$ is a structure that is homomorphically equivalent to $\mathbf{C}$, there exists a core $\mathbf{C}^{\prime}$ of $\mathbf{B}$ that is isomorphic to $\mathbf{C}$.
[…]
2.2 Query plans
For completeness and clarity, we precisely define the types of plans to be studied. For each $m\geq 0$, we define an $m$-suitable identification as an expression $j=k$ where $j,k\in[m]$. For a signature $\sigma$, a SPJU plan over $\sigma$ is inductively defined as follows: (basic) For each $R\in\sigma$, it holds that $R$ is a plan of arity $\mathrm{ar}(R)$. (select) When $p$ is a plan of arity $m$ and $\theta$ is a set of $m$-suitable identifications, $\sigma_{\theta}(p)$ is a plan of arity $m$. (project) When $p$ is a plan of arity $m$ and $j_{1},\ldots,j_{n}$ is a sequence of numbers from $[m]$, with $n\geq 0$, it holds that $\pi_{j_{1},\ldots,j_{n}}(p)$ is a plan of arity $n$. (join) When $p_{1},\ldots,p_{\ell}$ are plans of arity $m_{1},\ldots,m_{\ell}$, respectively, with $\ell\geq 1$, and $\theta$ is a set of $(m_{1}+\cdots+m_{\ell})$-suitable identifications, $\Join_{\theta}(p_{1},\ldots,p_{\ell})$ is a plan of arity $(m_{1}+\cdots+m_{\ell})$. (union) When $p_{1},p_{2}$ are plans of arity $m$, it holds that $p_{1}\cup p_{2}$ is a plan of arity $m$. In this paper, we will sometimes simply use the term plan to refer to an SPJU plan. We define an SPJ plan to be a SPJU plan that does not make use of union. Suppose that $Q$ is a relation of arity $m$. For any set $\theta$ of $m$-suitable identifications, we define $\sigma_{\theta}(Q)=\{(d_{1},\ldots,d_{m})\in Q\penalty 10000\ |\penalty 10000\ \mbox{for each $(j=k)\in\theta$, $d_{j}=d_{k}$ holds}\}.$ For any sequence $j_{1},\ldots,j_{n}\in[m]$, we define $\pi_{j_{1},\ldots,j_{n}}(Q)=\{(d_{j_{1}},\ldots,d_{j_{n}})\penalty 10000\ |\penalty 10000\ (d_{1},\ldots,d_{m})\in Q\}.$ Let $p$ be a SPJU plan over a signature $\sigma$, let $\mathbf{D}$ be a structure over $\sigma$, and let $m$ be the arity of $p$. We define the output of $p$ over $\mathbf{D}$, denoted by $\mathsf{out}(p,\mathbf{D})$, to be the arity $m$ relation over $D$ defined as follows. Note that, as is usual in this context, we understand that there is a unique tuple of arity 0, called the empty tuple. When $p$ has the form $R$, we define $\mathsf{out}(p,\mathbf{D})=R^{\mathbf{D}}.$ When $p$ has the form $\sigma_{\theta}(p^{\prime})$, we define $\mathsf{out}(p,\mathbf{D})=\sigma_{\theta}(\mathsf{out}(p^{\prime},\mathbf{D})).$ When $p$ has the form $\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, we define $\mathsf{out}(p,\mathbf{D})=\pi_{j_{1},\ldots,j_{n}}(\mathsf{out}(p^{\prime},\mathbf{D})).$ When $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, we define $\mathsf{out}(p,\mathbf{D})=\sigma_{\theta}(\mathsf{out}(p_{1},\mathbf{D})\times\cdots\times\mathsf{out}(p_{\ell},\mathbf{D})).$ When $p$ has the form $p_{1}\cup p_{2}$, we define $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p_{1},\mathbf{D})\cup\mathsf{out}(p_{2},\mathbf{D}).$ Observe that selection is a special case of join, in particular, it corresponds to a $1$-way join. We say that two plans $p,p^{\prime}$ over the same signature $\sigma$ are semantically equivalent when, for every structure $\mathbf{D}$ over $\sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{\prime},\mathbf{D})$. Further, for a set $\Sigma$ of keys over $\sigma$, we say that two plans $p,p^{\prime}$ over $\sigma$ are $\Sigma$-semantically equivalent when, for every structure $\mathbf{D}$ over $\sigma$ that satisfies $\Sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{out}(p^{\prime},\mathbf{D})$. A subplan of a plan is defined in the usual manner. Formally, we define a set $\mathsf{subplans}(p)$, for each plan $p$, inductively as follows: When $p$ has the form $R$, we define $\mathsf{subplans}(p)=\{p\}$. When $p$ has the form $\sigma_{\theta}(p^{\prime})$ or $\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, we define $\mathsf{subplans}(p)=\mathsf{subplans}(p^{\prime})\cup\{p\}.$ When $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, we define $\mathsf{subplans}(p)=(\bigcup_{i=1}^{\ell}\mathsf{subplans}(p_{i}))\cup\{p\}.$ When $p$ has the form $p_{1}\cup p_{2}$, we define $\mathsf{subplans}(p)=\mathsf{subplans}(p_{1})\cup\mathsf{subplans}(p_{2})\cup\{p\}.$ We then say that $q$ is a subplan of $p$ when $q\in\mathsf{subplans}(p)$.
2.3 Hypergraphs and tree decompositions
A hypergraph is a pair $(V,E)$ where $V$ is a set, called the vertex set, and $E\subseteq 2^{V}$ is a set called the edge set. A graph is a hypergraph where each edge has size $2$. When $H$ is a hypergraph, we use $V(H)$ and $E(H)$ to denote its vertex set and edge set, respectively. An isolated vertex of a hypergraph $H$ is a vertex not occurring in any edge, that is, an element of $V(H)\setminus\bigcup_{e\in E(H)}e$. When $H$ is a hypergraph and $S\subseteq V(H)$, we use $H[S]$ to denote the induced subhypergraph of $H$ on $S$, that is, the hypergraph with vertex set $S$ and with edge set $\{e\cap S\penalty 10000\ |\penalty 10000\ e\in E(H)\}$. When $\mathbf{A}$ is a structure over signature $\sigma$ we use $\mathbf{H}(\mathbf{A})$ to denote the hypergraph with vertex set $\mathsf{dom}(\mathbf{A})$ and edge set $\{\{\overline{a}\}\penalty 10000\ |\penalty 10000\ R\in\sigma,\overline{a}\in R^{\mathbf{A}}\}$. When $(\mathbf{A},\overline{a})$ is an open structure, we use $\mathbf{H}(\mathbf{A},\overline{a})$ to denote the hypergraph with vertex set $\mathsf{dom}(\mathbf{A})$ and with edge set $E(\mathbf{H}(\mathbf{A}))\cup\{\{\overline{a}\}\}$. A tree decomposition of a hypergraph $H$ is a pair $(T,\chi)$, where $T$ is a tree and $\chi$ is a labeling function $V(T)\rightarrow 2^{V(H)}$, i.e., it assigns each vertex of $T$ a subset of vertices of $H$, such that 1. (vertex coverage) for each $v\in V(H)$, there exists $t\in V(T)$ such that $v\in\chi(t)$, 2. (edge coverage) for each $e\in E(H)$, there exists $t\in V(T)$ such that $e\subseteq\chi(t)$, and 3. (connectivity) for each $v\in V(H)$, the set $\{t\in V(T)\penalty 10000\ |\penalty 10000\ v\in\chi(t)\}$ induces a connected subtree of $T$. Note that, if $H$ has no isolated vertices, condition $(1)$ is implied by condition $(2)$. A tree decomposition of an open structure $(\mathbf{A},\overline{a})$, is a tree decomposition of its hypergraph $\mathbf{H}(\mathbf{A},\overline{a})$.
2.4 Key constraints
A key constraint $\kappa$ over signature $\sigma$ is an expression $\textrm{key}(R)=K$, where $R/n\in\sigma$ and $K\subseteq[n]$. For an $n$-tuple $\overline{a}=(a_{1},\ldots,a_{n})$ and some $K=\{i_{1},\ldots,i_{m}\}\subseteq[n]$, for some $m\in[n]$, we write $\overline{a}[K]$ for the tuple $(a_{i_{1}},\ldots,a_{i_{m}})$. A structure $\mathbf{A}$ over signature $\sigma$ satisfies $\kappa$ if, for every two tuples $\overline{a},\overline{b}\in R^{\mathbf{A}}$, $\overline{a}[K]=\overline{b}[K]$ implies $\overline{a}=\overline{b}$. If, for some key $\kappa$ of the form $\textrm{key}(R)=K$, the set $K$ is a singleton, we say that $\kappa$ is a unary key. We say a structure satisfies a set of keys $\Sigma$, if it satisfies every key $\kappa\in\Sigma$. The chase is a well-known tool for enforcing key dependencies (see, e.g., [16, 8]). For an open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$ and a set $\Sigma$ of keys over $\sigma$ a chase step on $(\mathbf{A},\overline{a})$ w.r.t. $\Sigma$ is a substitution of the form $x\mapsto y$, which is applied globally on $\mathbf{A}$ and $\overline{a}$, where there are two tuples $\overline{b},\overline{b^{\prime}}\in R^{\mathbf{A}}$ and a key $\textrm{key}(R)=K$ for some $R/n\in\sigma$, such that $\overline{b}[K]=\overline{b^{\prime}}[K]$ but there exists some $i\in[n]\setminus K$ such that $\overline{b}[i]=x$ and $\overline{b^{\prime}}[i]=y$. The chase of $(\mathbf{A},\overline{a})$ w.r.t. $\Sigma$, denoted $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, is the result of repeated applications of chase steps until a fixpoint is reached; this fixpoint is known to be unique up to isomorphism. The chase of a structure $\mathbf{A}$ over signature $\sigma$ w.r.t. $\Sigma$, denoted $\textrm{Chase}_{\Sigma}(\mathbf{A})$, is defined analogously in the obvious way. We next define a measure for hypergraphs based on a notion, called color number, defined on conjunctive queries and introduced in [10]. For a structure $\mathbf{A}$ over signature $\sigma$ and a set $\Sigma$ of keys over $\sigma$, a valid coloring of $\mathbf{A}$ w.r.t. $\Sigma$ is a mapping $\mathsf{Col}_{\Sigma}^{\mathbf{A}}:\mathsf{dom}(\mathbf{A})\rightarrow 2^{\{1,\ldots,k\}}$, for some $k>0$, which assigns to each element of the universe of $\mathbf{A}$ a subset of $k$ colors such that the following hold: 1. For each key $\textrm{key}(R)=\{i_{1},\ldots,i_{m}\}\in\Sigma$, i.e., $R/n\in\sigma$ and $\{i_{1},\ldots,i_{m}\}\subseteq[n]$, we have that for every $\overline{a}\in R^{\mathbf{A}}$ and every $i\in[n]\setminus\{i_{1},\ldots,i_{m}\}$ it holds that $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(\overline{a}[i])\subseteq\bigcup_{j\in[m]}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(\overline{a}[i_{j}])$. 2. There exists an element $x\in\mathsf{dom}(\mathbf{A})$ such that $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)\neq\emptyset$. For a structure $\mathbf{A}$ over signature $\sigma$, a set $S\subseteq\mathsf{dom}(\mathbf{A})$, and a set $\Sigma$ of keys over $\sigma$ the color number of $S$ w.r.t. $\mathbf{A}$ and $\Sigma$, denoted $C_{\Sigma}^{\mathbf{A}}(S)$, is defined as the maximal ratio of colors assigned to the elements in $S$ over the maximum number of colors assigned to the elements of some tuple in $\mathbf{A}$ which is attainable by any valid coloring (for any $k$) of $\mathbf{A}$ w.r.t. $\Sigma$. Formally, where $\max_{\mathsf{Col}_{\Sigma}^{\mathbf{A}}}$ ranges over all such valid colorings: $C_{\Sigma}^{\mathbf{A}}(S)=\max_{\mathsf{Col}_{\Sigma}^{\mathbf{A}}}\frac{|\bigcup_{x\in S}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)|}{\max_{\overline{b}\in\mathbf{A}}|\bigcup_{x\in\overline{b}}\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)|}.$ A crucial observation is that for a coloring that maximises the color number of $S$, we have that $x\notin S$ implies $\mathsf{Col}_{\Sigma}^{\mathbf{A}}(x)=\emptyset$. Note that, for a structure $\mathbf{A}$ over signature $\sigma$, the color number $C_{\Sigma}^{\mathbf{A}}(S)$, as defined here, is equal to the color number, as defined in [10], of the conjunctive query whose atoms are those in $\{R(v_{1},\ldots,v_{k})\penalty 10000\ |\penalty 10000\ R\in\sigma,(v_{1},\ldots,v_{k})\in R\}$, and whose output variables are $S$. We refer the reader to [10] for more information on the color number, but note the following fact here.
Proposition 2.3.
(refer to [10, Section 3]) Let $\mathbf{B}$ be a structure, and let $S$ be a subset of $\mathsf{dom}(\mathbf{B})$. The value $C^{\mathbf{B}}_{\emptyset}(S)$ is the fractional edge cover number of the induced hypergraph of $\mathbf{H}(\mathbf{B})$ on $S$.
A fractional edge cover of a hypergraph $G=(V,F)$ is a mapping $x:F\to[0,\infty)$ such that for each $v\in V$, it holds that $\sum_{e\in E,v\in e}x(e)\geq 1$; the sum $\sum_{e\in E}x(e)$ is the weight of $x$. The fractional edge cover number of $G$ is the minimum weight over all fractional edge covers of $G$; it follows from the theory of linear programming that a minimum exists [12]. Let us view $C_{\Sigma}^{\mathbf{A}}$ as a mapping from the power set of $\mathsf{dom}(\mathbf{A})$ to $\mathbb{Q}^{+}$. We define the following width notions. For an open structure $(\mathbf{A},\overline{a})$ over signature $\sigma$, a set $\Sigma$ of keys over $\sigma$, and a tree decomposition $F=(T,\chi)$ of $\mathbf{H}(\mathbf{A},\overline{a})$, let $C_{\Sigma}^{\mathbf{A}}\textup{-width}(F)=\max\{C_{\Sigma}^{\mathbf{A}}(\chi(t))\penalty 10000\ |\penalty 10000\ t\in V(T)\}$. Further, let $C_{\Sigma}^{\mathbf{A}}\textup{-width}(\mathbf{A},\overline{a})=\min_{F}C_{\Sigma}^{\mathbf{A}}\textup{-width}(F)$, where $F$ ranges over all tree decompositions of $\mathbf{H}(\mathbf{A},\overline{a})$.
3 Degree Notions and Main Theorem Statement
3.1 Degree notions
A key measure in the context of query optimization is the size of the output of a query. Here, we are interested in an asymptotic measure: the output degree of a query plan $p$, which, informally, is the smallest possible degree of a polynomial bounding the output size of $p$ on any structure $\mathbf{D}$ satisfying given constraints, as a function of the maximum relation size $M_{\mathbf{D}}$; the formal definition is the following.
Definition 3.1 (Output Degree).
Let $\Sigma$ be a set of keys over signature $\sigma$. Let $p$ be a plan over $\sigma$, and let $d\in\mathbb{Q}$. The plan $p$ has output degree $\leq d$ if there exists a function $h:\mathbb{N}\to\mathbb{N}$, with $h(M)\in O(M^{d})$, such that for every structure $\mathbf{D}$ satisfying $\Sigma$, it holds that $|\mathsf{out}(p,\mathbf{D})|\leq h(M_{\mathbf{D}})$. We say that $p$ has output degree $d$ if it has output degree $\leq d$, but for all $\epsilon>0$, it does not have output degree $\leq(d-\epsilon)$.
Instead of just focusing on some plan as a whole, we are interested in bounding also the size of the output of every subplan of the original plan, or in other words, bounding the size of intermediate results obtained during the execution of the plan. To do so, we say a plan $p$ has intermediate degree $d$ if there exists a degree $d$ polynomial bounding the sizes of all intermediate relations computed during the execution of $p$ on any structure $\mathbf{D}$ satisfying given constraints, as a function of $M_{\mathbf{D}}$—and again, there is no lower degree polynomial providing such a bound.
Definition 3.2 (Intermediate Degree).
Let $\Sigma$ be a set of keys over signature $\sigma$. Let $p$ be a plan over $\sigma$, and let $d\in\mathbb{Q}$. The plan $p$ has intermediate degree $\leq d$ if there exists a function $h:\mathbb{N}\to\mathbb{N}$, with $h(M)\in O(M^{d})$, such that for every structure $\mathbf{D}$ satisfying $\Sigma$, and for every subplan $q$ of $p$, it holds that $|\mathsf{out}(q,\mathbf{D})|\leq h(M_{\mathbf{D}})$. We say that $p$ has intermediate degree $d$ if it has intermediate degree $\leq d$, but for all $\epsilon>0$, it does not have intermediate degree $\leq(d-\epsilon)$.
Note that the above definitions of output degree and intermediate degree are with respect to the keys $\Sigma$, but we omit $\Sigma$ when it is clear from context for the sake of readability. Let $\Sigma$ be a set of keys. We say that a plan $p$ has the SPJ-best possible intermediate degree if there exists $d\in\mathbb{Q}$ such that $p$ has intermediate degree $d$, and for every SPJ plan $p^{\prime}$ that is $\Sigma$-semantically equivalent to $p$, there is no value $\epsilon>0$ such that the plan $p^{\prime}$ has intermediate degree $\leq(d-\epsilon)$. In other words, no $\Sigma$-semantically equivalent SPJ plan achieves an intermediate degree that is strictly lower than that of $p$. Analogously, we say that a plan $p$ has the SPJU-best possible intermediate degree if there exists $d\in\mathbb{Q}$ such that $p$ has intermediate degree $d$, and for every SPJU plan $p^{\prime}$ that is $\Sigma$-semantically equivalent to $p$, there is no value $\epsilon>0$ such that the plan $p^{\prime}$ has intermediate degree $\leq(d-\epsilon)$. In the appendix, we give a number of examples of plans along with computations of the output degree and the best possible intermediate degree.
3.2 Main theorem statement
We are now ready to formulate our main theorem.
Theorem 3.3 (Main theorem).
There exists an exponential-time algorithm that, given a signature $\sigma$, a set $\Sigma$ of unary keys over $\sigma$, and an SPJU plan $p$ over $\sigma$, outputs a SPJU plan $p^{\prime}$ satisfying the following properties: $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$, $p^{\prime}$ has the SPJU-best possible intermediate degree.
By an exponential-time algorithm, we mean an algorithm that runs within time $O(2^{Q(n)})$, where $n$ is the input size and $Q$ is a polynomial. The principal argumentation needed to establish our main theorem is located in the case where this theorem is specialized to SPJ queries. In this case, which we present next, we can describe the intermediate degree of the output plan quite directly, but require the following notion to do so. A $p$-representation of a plan $p$ is an open structure $(\mathbf{A},\overline{a})$ where, for any structure $\mathbf{D}$, we have $\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$. A formal treatment of $p$-representation and related notions utilized to show our results will follow in Subsection 4.1.
Theorem 3.4.
There exists an exponential-time algorithm that, given a signature $\sigma$, a set $\Sigma$ of unary keys over $\sigma$, and an SPJ plan $p$ over $\sigma$, outputs a SPJ plan $p^{\prime}$ satisfying the following properties: $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$, $p^{\prime}$ has the SPJ-best possible intermediate degree. Moreover, suppose that $p$ is an SPJ plan over $\sigma$ and that $(\mathbf{A},\overline{a})$ is a $p$-representation; let $(\mathbf{C},\overline{c})$ be a core of $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, and let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. Then, when the algorithm is given $p$ as input, the output plan $p^{\prime}$ has intermediate degree $d$, and also has size polynomial in the size of $p$.
Some remarks are in order. First, note that in Theorem 3.4, the intermediate degree of each plan $p^{\prime}$ can be described quite explicitly. Note also that given an SPJ plan $p$, the intermediate degree $d$ of the plan $p^{\prime}$ given by the algorithm of Theorem 3.4 is computable in exponential time from $p$; this follows directly from the exponential-time computability of $(\mathbf{C},\overline{c})$ from $p$ and the polynomial time computability of the color number, which follows from results in Sections 3 and 4 from [10]. In turn, given an SPJU plan $p$, the intermediate degree of a plan $p^{\prime}$ output by Theorem 3.3’s algorithm is also computable in exponential time from $p$, which follows from the above fact and the construction of $p^{\prime}$ in the proof of Theorem 3.3. We also wish to remark that, while our main theorem only considers constraints which are unary keys, this limitation arises from the literature on output degree bounds; extending the literature’s bounds is likely to be challenging, for this would constitute significant progress towards a long-standing open problem in information theory [10, Section 1].
3.3 Algorithmic aspects
We here discuss algorithms for evaluating well-behaved plans in time that is close to a plan’s intermediate degree; this issue was motivated in the introduction. We define a plan $p$ to be well-behaved if for each subplan $q$ that is a join $\Join_{\theta}(p_{1},\ldots,p_{\ell})$ of plans $p_{1},\ldots,p_{\ell}$ having arities $m_{1},\ldots,m_{\ell}$, the following is satisfied, when we let $s$ denote the sum $m_{1}+\cdots+m_{\ell}$: there exists an index $k\in[s]$ such that for each index $i\in[s]\setminus([m_{1}]\cup\{k\})$, there exists an index $j\in[m_{1}]\cup\{k\}$ such that $(i=j)\in\theta$. In words, each join adds at most one column of new information to the plan $p_{1}$: letting $J$ denote the set containing (1) the indices in $m_{1}]$, which correspond to $p_{1}$, and (2) another chosen index $k$; it must be that every index $i\notin J$ is equated with an index in $J$. This property allows for the efficient computation of such joins and, thus, well-behaved plans; we will make this formal in Proposition 3.6. First, we point out that our minimization algorithms can be implemented in such a way that they always output well-behaved plans.
Theorem 3.5.
The algorithms of Theorems 3.3 and 3.4 can be implemented in such a way so that each plan $p^{\prime}$ output by them is well-behaved.
This easily follows from the fact that the algorithm of Theorem 4.7 always outputs well-behaved SPJ plans and that the SPJU plan output by the algorithm of Theorem 3.3 is simply the union of well-behaved SPJ plans, which is also a well-behaved plan. The following proposition shows that, given a well-behaved plan, apart from a multiplicative overhead depending on the plan, the time needed to algorithmically evaluate the plan on a database $\mathbf{D}$ is at most $|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d}$, where $d$ is the intermediate degree of the plan—so in essence, the intermediate degree of the plan dictates the polynomial degree of evaluation (as a function of $M_{\mathbf{D}}$). To argue this, we in essence use the argumentation for [2, Theorem 6], and we assume a standard random access model of computation, with a uniform cost measure.
Proposition 3.6.
There exists an algorithm that, given a well-behaved SPJU plan $p$ of intermediate degree $\leq d$ and a structure $\mathbf{D}$, evaluates $p$ on $\mathbf{D}$ in time $O(|p|^{2}*|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d})$.
To argue this proposition, we consider an algorithm that evaluates the plan $p$ inductively. It suffices to show that each subplan $q\in\mathsf{subplans}(p)$ can be computed in time $O(|p|*|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d})$. By the definition of intermediate degree, each subplan $q$ has at most $O(M_{\mathbf{D}}^{d})$ answers (tuples in $\mathsf{out}(p,\mathbf{D})$). When $q$ is a basic plan, the required bound holds immediately. When $q$ is a select or project subplan, i.e., of the form $\sigma_{\theta}(q^{\prime})$ or $\pi_{j_{1},\ldots,j_{n}}(q^{\prime})$, respectively, the answers of $q$ can be naturally derived from the answers of the immediate subplan $q^{\prime}$, within the given time bound. When $q$ is a union subplan, i.e., of the form $q_{1}\cup q_{2}$, it is immediate that the answers of $q$ can be derived by combining the answers of $q_{1}$ and $q_{2}$ and eliminating duplicates, which is feasible in the given time bound. The key point is that, when $q$ is a join subplan, by well-behavedness, it is of the form $\Join_{\theta}(q_{1},\ldots,q_{\ell})$ and $\mathsf{out}(q,\mathbf{D})$ can be computed by maintaining a relation that is a subset of $\mathsf{out}(q_{1},\mathbf{D})\times\mathsf{dom}(\mathbf{D})$. When evaluating the multiway join, we can thus evaluate the sequence of binary joins $\Join_{\theta_{\ell-1}}(\ldots(\Join_{\theta_{2}}(\Join_{\theta_{1}}(q_{1},q_{2}),q_{3}),\ldots),q_{\ell})$, where $\theta_{1},\ldots,\theta_{\ell-1}$ are the corresponding equalities from $\theta$ for the join partners in question. Each binary join can be evaluated in time $O(|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d})$ by [9], and in total, there are at most $|p|$ many joins, giving us the desired runtime of $O(|p|*|\mathsf{dom}(\mathbf{D})|*M_{\mathbf{D}}^{d})$ to evaluate the subplan $q$.
4 SPJ Plans
This section is dedicated to proving Theorem 3.4, which, apart from being of interest in its own right, serves as a building block towards proving our main result in the next section. Proving Theorem 3.4 essentially requires us to show two statements: (i) we can synthesize a plan, which is $\Sigma$-equivalent to the original one and has intermediate degree $\leq d$ and (ii) any plan which is $\Sigma$-equivalent to the original one has a subplan with output degree at least $d$. We will formalize and prove (i) and (ii) in Theorems 4.7 and 4.8, respectively, in Subsection 4.2. To show the theorem, we make two more basic observations on the introduced degree notions. We begin by presenting a definition that will aid us in arguing that a plan has a certain output degree. We say that $p$ has output degree $\geq d$ if there exists a function $g:\mathbb{N}\to\mathbb{N}$ with $g(M)\in\Omega(M^{d})$ and there exists an infinite sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures with $M_{\mathbf{D}_{n}}$ unbounded such that (for all $n\geq 1$) $|\mathsf{out}(p,\mathbf{D}_{n})|\geq g(M_{\mathbf{D}_{n}})$. We then have the following lemma.
Lemma 4.1.
Let $d\geq 0$. If a plan has output degree $\geq d$, then for any $\epsilon>0$, it does not have output degree $\leq(d-\epsilon)$.
Next, we characterize the intermediate degree of a plan in terms of the output degrees of subplans, in the following sense.
Lemma 4.2.
Let $d\geq 0$. A plan $p$ has intermediate degree $\leq d$ if and only if each subplan $q$ of $p$ has output degree $\leq d$.
Theorem 3.4 can then be argued as follows: Given a plan $p$, we first compute the open structure $(\mathbf{A},\overline{a})$, which can be done in polynomial time by Theorem 4.6. Then, the algorithm of Theorem 3.4 computes $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, again in polynomial time. It then computes the core of this open structure; this can be performed in exponential time. Let $(\mathbf{C},\overline{c})$ be this core. Note that the size of $(\mathbf{C},\overline{c})$ is bounded above by the size of $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})$, since the core of a structure is a substructure of the structure. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. From $(\mathbf{C},\overline{c})$, we can then synthesize the desired plan $p^{\prime}$ of intermediate degree $\leq d$ by statement (i), formalized in Theorem 4.7, from above. Consider any plan $p^{\prime\prime}$ that is $\Sigma$-semantically equivalent to $p^{\prime}$. By statement (ii), formalized in Theorem 4.8, $p^{\prime\prime}$ has a subplan with output degree $\geq d$. For any $\epsilon>0$, by Lemma 4.1, this subplan does not have output degree $\leq(d-\epsilon)$, and so by Lemma 4.2, $p^{\prime\prime}$ does not have intermediate degree $\leq(d-\epsilon)$. Hence, we have computed a $\Sigma$-equivalent plan $p^{\prime}$ of polynomial size and of intermediate degree $d$, which is the best possible intermediate degree. In the following, we will formalize the connection from plans to open structures and auxiliary notions, which are needed to prove the theorems formalizing statements (i) and (ii) in Subsection 4.2.
4.1 Relating Queries to Structures
There is a known correspondence between Boolean SPJ plans and structures [5]; by a Boolean SPJ plan, we mean one that evaluates to true or false (or alternatively, a relation of arity 0) on a given structure $\mathbf{D}$. This correspondence yields that each such plan can be translated to a structure $\mathbf{A}$ such that, for an arbitrary structure $\mathbf{D}$, the plan is true on $\mathbf{D}$ if and only if there is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$. We prove a version of this correspondence in Theorem 4.6 for SPJ plans that can output relations of any finite arity, and not just relations of arity 0. In order to formulate our version, we utilize the notion of open structure.
Definition 4.3 ($p$-Representation).
Let $p$ be an SPJ plan over signature $\sigma$. A $p$-representation is an open structure $(\mathbf{A},\overline{a})$ over $\sigma$, where for each structure $\mathbf{D}$ over $\sigma$, it holds that $\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$.
Apart from showing that any SPJ plan $p$ can be translated to a $p$-representation, we also want to provide a tree decomposition of this open structure with certain additional properties, which we call $p$-decomposition. This allows us to relate plans to structural, decomposition-based measures. While the translation that we present is, on a high level, along the lines of known translations from SPJ plans to conjunctive queries, we emphasize that our translation also provides a tree decomposition that will allow us to analyze the intermediate degree. In particular, for a given plan, each subplan will correspond to a vertex of the tree decomposition whose associated set of elements corresponds to the free variables of the subplan. We say a tree decomposition $(T,\chi)$ is rooted if there is a distinguished node $v_{0}\in V(T)$, called the root of $T$. The formal definition of $p$-decomposition follows.
Definition 4.4 ($p$-Decomposition).
Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ a $p$-representation. A $p$-decomposition of $(\mathbf{A},\overline{a})$ is a rooted tree decomposition $E=(T,\chi)$ of $(\mathbf{A},\overline{a})$ such that there exists a bijection $\alpha:\mathsf{subplans}(p)\to V(T)$, where $\alpha(p)$ is the root of $T$, and a map $\beta:\mathsf{subplans}(p)\to\mathsf{dom}(\mathbf{A})^{*}$, where $\beta(p)$ is equal to $\overline{a}$, where for each $q\in\mathsf{subplans}(p)$, we have that $\{\beta(q)\}=\chi(\alpha(q))$.
In the definition of $p$-decomposition, each subplan $q$ of $p$ corresponds to a vertex of the tree $T$, via $\alpha$, and for each such subplan $q$, it holds that $\beta(q)$ is a tuple that lists, in some order (and possibly with repetition), the elements of the set to which $\chi$ maps q’s tree vertex. Let $p$ be an SPJ plan over signature $\sigma$, $(\mathbf{A},\overline{a})$ be a $p$-representation, and $E=(T,\chi)$ be a rooted $p$-decomposition $E=(T,\chi)$ of $(\mathbf{A},\overline{a})$ with associated mappings $\alpha$ and $\beta$. For a vertex $v\in T$, let $S_{v}$ be the union of $\chi(u)$ over all descendants $u$ of $v$ (including $v$ itself) and let $\mathbf{A}_{\leq v,E}$ denote the induced substructure $\mathbf{A}[S_{v}]$. We say a subplan $q\in\mathsf{subplans}(p)$ satisfies the containment property w.r.t. a $p$-decomposition if, for each structure $\mathbf{D}$ over $\sigma$, the relationship $\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}_{\leq\alpha(q),E},\beta(q),\mathbf{D})$ holds. Further, we say a $p$-decomposition satisfies the containment property, if every subplan $q\in\mathsf{subplans}(p)$ satisfies the containment property w.r.t. the $p$-decomposition. The containment property will allow us to lower bound the sizes of intermediate relations of $p$ using a $p$-representation along with a $p$-decomposition satisfying the containment property.
Example 4.5.
Consider the signature $\sigma$ containing the binary relations $R/2$ and $S/2$ and the SPJ plan $p=\Join_{(1=3)}(R,\pi_{1}(S))$ over $\sigma$. Then it can be readily seen, with structure $\mathbf{A}=\{R(a,b),S(a,d)\}$ and tuple $\overline{a}=(a,b,a)$, that $(\mathbf{A},\overline{a})$ is a $p$-representation. Consider the following trees: [Uncaptioned image] On the left, we have the plan $p$ represented as a tree (each node represents the subplan of its subtree), while on the right we have some tree $T$. We provide mappings $\alpha$, $\beta$, and $\chi$ such that $(T,\chi)$ is a $p$-decomposition via mappings $\alpha$ and $\beta$. First, $\alpha$ is the obvious bijection from the subplans of $p$ to $V(T)$ with $\alpha(p)=v_{0}$, $\alpha(R)=v_{1}$, $\alpha(\pi_{1}(S))=v_{2}$, and $\alpha(S)=v_{3}$. Further, we define $\beta$ as $\beta(p)=\overline{a}$, $\beta(R)=(a,b)$, $\beta(\pi_{1}(S))=(a)$, and $\beta(S)=(a,d)$. Lastly, define $\chi(\alpha(q))$ as $\{\beta(q)\}$ for all subplans $q$ of $p$. It can be readily verified that $(T,\chi)$ is indeed a $p$-decomposition, and that it satisfies the containment property.
We show with the following theorem, that a $p$-representation and a $p$-decomposition satisfying the containment property always exist and are efficiently computable, via an inductive argument over the structure of $p$ given in the appendix.
Theorem 4.6.
There exists a polynomial-time algorithm that, given an SPJ plan $p$ over signature $\sigma$, outputs a $p$-representation over $\sigma$ along with a $p$-decomposition of this $p$-representation which satisfies the containment property.
4.2 Intermediate Size Bounds
We devote this subsection to formalizing and establishing the statements used to prove Theorem 3.4. We start with statement (i), i.e. we show that we can synthesize a plan which is $\Sigma$-equivalent to the one given originally and which has the desired intermediate degree.
Theorem 4.7 (Plan synthesis).
There exists an exponential-time algorithm that, given a set $\Sigma$ of unary keys over a signature $\sigma$ and an open structure $(\mathbf{C},\overline{c})$ over $\sigma$ with $\textrm{Chase}_{\Sigma}(\mathbf{C},\overline{c})=(\mathbf{C},\overline{c})$ and which is a $p$-representation for some plan $p$, outputs a well-behaved plan $p^{\prime}$ where (a) $p^{\prime}$ is $\Sigma$-semantically equivalent to $p$, and (b) $p^{\prime}$ has intermediate degree $\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$.
To show the above theorem, we combine proof techniques from [2] and [10]. As a first step, following arguments from Section 4 of [10], we transform the open structure $(\mathbf{C},\overline{c})$ together with $\Sigma$ to an open structure $(\mathbf{C}^{\prime},\overline{c})$ (over a modified signature $\sigma^{\prime}$) which has no constraints posed over it. This transformation can be achieved via the application of a well-behaved SPJ plan of intermediate degree 1, and where every new relation is of a size that is bounded from above by an original relation. Crucially, for any structure $\mathbf{D}$ satisfying $\Sigma$, we have that $(\mathbf{C},\overline{c})$ returns the same answers on $\mathbf{D}$ as $(\mathbf{C}^{\prime},\overline{c})$ on an analogous transformation of $\mathbf{D}$. Moreover, we have that $C_{\emptyset}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$ and it thus suffices to show the result for the empty set of unary keys $\Sigma^{\prime}=\emptyset$ along with $(\mathbf{C}^{\prime},\overline{c})$, in place of $\Sigma$ with $(\mathbf{C},\overline{c})$. Finally, we compute a tree decomposition $F$ of $(\mathbf{C}^{\prime},\overline{c})$ where $C_{\Sigma^{\prime}}^{\mathbf{C}^{\prime}}\textup{-width}(F)$ is equal to $C_{\Sigma^{\prime}}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})$, which can be done in exponential time, and then compute a SPJ plan $p^{\prime}$ following the tree composition using standard techniques from [2]. We provide a detailed proof in the appendix. We continue to show that statement (ii) also holds. To that end, let us first formalize the statement. In essence, it provides a lower bound on the intermediate degree of any plan that is $\Sigma$-semantically equivalent to $p$.
Theorem 4.8.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ be a $p$-representation. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$, and let $(\mathbf{C},\overline{c})$ be a core of $(\mathbf{B},\overline{b})$. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. For any SPJ plan $p^{+}$ that is $\Sigma$-semantically equivalent to $p$, it holds that $p^{+}$ has a subplan with output degree $\geq d$.
Before proceeding with the proof, we establish two preparatory lemmas that will aid us in showing the lower bound on the intermediate degree. First, we show that for an open structure $(\mathbf{C},\overline{c})$ which is a core of the chase of a $p$-representation and which has width $d$, there is a vertex in any tree decomposition of $(\mathbf{C},\overline{c})$ witnessing that the intermediate results grow in accordance with a degree $d$ polynomial.
Lemma 4.9.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ be a $p$-representation. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$, and let $(\mathbf{C},\overline{c})$ be a core of $(\mathbf{B},\overline{b})$. Let $d=C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$. For any tree decomposition $(T,\chi)$ of $(\mathbf{C},\overline{c})$, there exists a vertex $t_{0}$ of $T$, a constant $c^{\prime}>0$, and a sequence of structures $(\mathbf{D}_{n})_{n\geq 1}$ with $M_{\mathbf{D}_{n}}$ unbounded where $|\mathsf{homs}(\mathbf{C},\chi(t_{0}),\mathbf{D}_{n})|\geq c^{\prime}M_{\mathbf{D}_{n}}^{d}$.
We next give a lemma whose assumptions are a relaxation of (that is, they are implied by) the condition that $\mathbf{C}$ is a core of $\mathbf{B}$. It allows us to lower bound, for any subset $S\subseteq\mathsf{dom}(\mathbf{B})$, the size of $\mathsf{homs}(\mathbf{B},S,\cdot)$ by the size of $\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\cdot)$.
Lemma 4.10.
Suppose that $\mathbf{B},\mathbf{C}$ are structures over signature $\sigma$ such that $\mathbf{C}$ is a substructure of $\mathbf{B}$, and there exists a retraction $h$ from $\mathbf{B}$ to $\mathbf{C}$. Then, for any subset $S\subseteq\mathsf{dom}(\mathbf{B})$ and any structure $\mathbf{D}$ over $\sigma$, it holds that $|\mathsf{homs}(\mathbf{B},S,\mathbf{D})|\geq|\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\mathbf{D})|$.
We proceed to prove the final statement needed to show Theorem 3.4.
Proof 4.11 (Proof of Theorem 4.8).
[…]
5 SPJU Plans
Building on the established understanding of intermediate degree for SPJ plans from the previous section, we are now ready to prove Theorem 3.3. It is well-known that each SPJU plan $p$ is $\Sigma$-semantically equivalent to some plan $p^{\prime}$ which is a union $p_{1}\cup\cdots\cup p_{m}$ of SPJ plans. Let us note that the output degree of an SPJU plan can then be characterized as follows: for such a $\Sigma$-semantically equivalent plan $p^{\prime}$, the output degree of $p$ is equal to the maximum output degree over the plans $p_{1},\ldots,p_{m}$. To argue about the intermediate degree of an SPJU plan, we first give a theorem showing that a natural such translation to a union of SPJ plans exists and can be computed in exponential time, which has the additional property that the relation computed by a subplan of one of the SPJ plans is a subset of some relation computed by a subplan of the original SPJU plan; this inclusion then facilitates proving lower bounds on the intermediate degree of the original SPJU plan.
Theorem 5.1.
There exists an exponential-time algorithm that, given an SPJU plan $p$, outputs a sequence of SPJ plans $p_{1},\ldots,p_{m}$, a sequence of maps $\alpha_{1},\ldots,\alpha_{m}$, and a sequence of open structures $(\mathbf{A}_{1},\overline{a_{1}}),\ldots,(\mathbf{A}_{m},\overline{a_{m}})$ such that the following holds: 1. $p$ is semantically equivalent to $p_{1}\cup\cdots\cup p_{m}$, 2. for each $i=1,\ldots,m$, we have that $\alpha_{i}:\mathsf{subplans}(p_{i})\to\mathsf{subplans}(p)$ is a map such that (for each structure $\mathbf{D}$) when $q\in\mathsf{subplans}(p_{i})$, the containment $\mathsf{out}(q,\mathbf{D})\subseteq\mathsf{out}(\alpha_{i}(q),\mathbf{D})$ holds, 3. for each $i=1,\ldots,m$, the open structure $(\mathbf{A}_{i},\overline{a_{i}})$ is a $p_{i}$-representation, and $i\neq j$ implies that there is no homomorphism from $(\mathbf{A}_{i},\overline{a_{i}})$ to $(\mathbf{A}_{j},\overline{a_{j}})$.
On a high level, Theorem 3.3 can then be proved as follows. For some SPJU plan $p$, the algorithm first computes this convenient set of SPJ plans $p_{1},\ldots,p_{m}$ via Theorem 5.1. Then, for each obtained plan $p_{i}$, it computes a core $(\mathbf{C}_{i},\overline{c_{i}})$ of the chase of a $p$-representation, and then applies Theorem 3.4 to obtain a value $d_{i}$, and a plan $p^{\prime}_{i}$ such that $p_{i}$, $d_{i}$, and $p^{\prime}_{i}$ satisfy the properties given in that theorem statement. The algorithm outputs the value $d=\max_{i\in[m]}d_{i}$ and the plan $p^{\prime}=p^{\prime}_{1}\cup\cdots\cup p^{\prime}_{m}$. Since this algorithm outputs a union of well-behaved SPJ plans, also the resulting SPJU plan is well-behaved. It is easy to see that $p^{\prime}$ has intermediate degree $d$, it just remains to show that $d$ is indeed the SPJU-best possible intermediate degree. In essence, we prove that for any SPJU $q^{\prime}$ that is semantically equivalent to $p^{\prime}$, we have that from the sequence of SPJ plans $q^{\prime}_{1},\ldots,q^{\prime}_{n}$ obtained by applying Theorem 5.1 to $q^{\prime}$, there exists an index $\ell\in[n]$ such that $q^{\prime}_{\ell}$ does not have (for any $\epsilon>0$) intermediate degree $\leq(d-\epsilon)$ implying by condition (2) of Theorem 5.1 that $q^{\prime}$ does not have (for any $\epsilon>0$) intermediate degree $\leq(d-\epsilon)$. Full details can be found in the appendix.
6 Conclusion
The intermediate degree asymptotically characterizes the size of intermediate relations obtained during the execution of a query plan. For SPJU plans, we have shown how to optimize for this measure among all semantically equivalent SPJU plans with respect to a given set of unary key constraints. A natural way to extend our work is to consider either more expressive query languages, by adding more operators, or more expressive constraints, such as arbitrary functional dependencies.
Appendix A Examples
A.1 Example Illustrating Structure-related Notions (cf. Section 2)
Let $\sigma$ be the signature containing a single relation $E$ of arity $2$, and let $\mathbf{A}_{0}$ be the structure with universe $\mathsf{dom}(\mathbf{A}_{0})=\{u,v_{1},v_{2}\}$ and with relation $E^{\mathbf{A}_{0}}=\{(u,v_{1}),(u,v_{2})\}$. Consider the open structure $(\mathbf{A}_{0},(u))$; when $\mathbf{D}$ is a directed graph, we have that $\mathsf{homs}(\mathbf{A}_{0},(u),\mathbf{D})$ is equal to the arity $1$ relation containing each vertex (of $\mathbf{D}$) having an outgoing edge. Next, consider the open structure $(\mathbf{A}_{0},(v_{1},v_{2}))$; when $\mathbf{D}$ is a directed graph, we have that $\mathsf{homs}(\mathbf{A}_{0},(v_{1},v_{2}),\mathbf{D})$ is equal to the arity $2$ relation containing each pair of vertices (of $\mathbf{D}$) that each receive an incoming edge from a common vertex. This relation is symmetric, and when $x$ is a vertex of $\mathbf{D}$, this relation contains the pair $(x,x)$ if and only if $x$ receives an incoming edge (from some vertex). Further, for each $i=1,2$, let $\mathbf{A}_{i}$ be the structure with universe $\mathsf{dom}(\mathbf{A}_{i})=\{u,v_{i}\}$ and with relation $E^{\mathbf{A}_{i}}=\{(u,v_{i})\}$. Clearly, each of $\mathbf{A}_{1},\mathbf{A}_{2}$ is a substructure of $\mathbf{A}_{0}$. In addition, we have that the map fixing $u$ and $v_{1}$ and sending $v_{2}$ to $v_{1}$ is a retraction from $\mathbf{A}_{0}$ to $\mathbf{A}_{1}$; dually, the map fixing $u$ and $v_{2}$ and sending $v_{1}$ to $v_{2}$ is a retraction from $\mathbf{A}_{0}$ to $\mathbf{A}_{2}$. Each of $\mathbf{A}_{1},\mathbf{A}_{2}$ is a core of $\mathbf{A}_{0}$ (and of itself): to argue this, we need to argue that (say) $\mathbf{A}_{1}$ has no retraction to a proper substructure $\mathbf{A}^{\prime}$ (of $\mathbf{A}_{1}$). (With respect to $\mathbf{A}_{0}$, the structures $\mathbf{A}_{1}$, $\mathbf{A}_{2}$ are symmetric to each other.) We argue this as follows: if we had such a proper substructure, it would have a size $1$ universe, and the retraction would map each element in $\mathsf{dom}(\mathbf{A}_{1})$ to the single element $x$ in that size $1$ universe. But then this proper substructure $\mathbf{A}^{\prime}$ would have $E^{\mathbf{A}^{\prime}}=\{(x,x)\}$, which could not be a subset of $E^{\mathbf{A}_{1}}$, contradicting that $\mathbf{A}^{\prime}$ is a substructure of $\mathbf{A}_{1}$.
A.2 Examples of plans and computations of degree notions (cf. Section 3)
It follows from previous work that the output degree of an SPJ plan can be computed as follows.
Proposition A.1 (follows from [10]).
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $p$ be an SPJ plan over $\sigma$ and $(\mathbf{A},\overline{a})$ be a $p$-representation. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$. Then, $p$ has output degree $C^{\mathbf{B}}_{\Sigma}(\{\overline{b}\})$.
We remark that, as a consequence of this proposition, the output degree of a given SPJU plan can also be computed by converting the plan into a semantically equivalent plan that is a union of SPJ plans and taking the maximum output degree over all these SPJ plans. In the following, we give a number of examples of plans along with computations of the the output degree and the best possible intermediate degree. Throughout, we let $E$ be a relation symbol of arity $2$, and we let $\mathbf{A}$ be the structure with domain $\{a,b,c\}$ and the relation $E^{\mathbf{A}}=\{(a,b),(b,c),(c,a),(c,c)\}$. We set $\Sigma$ to be the empty set. We will use $\epsilon$ to denote the empty tuple.
Example A.2.
(output degree 3/2, best possible intermediate degree 3/2) Consider the plan $p_{1}=\pi_{1,2,4}\Join_{\theta}(E,E,E,E)$ where $\theta=\{1=6,2=3,4=5,5=7,7=8\}$. The definition of this set $\theta$ is based on the tuple $(a,b,b,c,c,a,c,c)$, consisting of domain elements of $\mathbf{A}$. We have that $(\mathbf{A},(a,b,c))$ is a $p_{1}$-representation. By Proposition A.1, the output degree of $p_{1}$ is $C^{A}_{\Sigma}(\{a,b,c\})$, which by Proposition 2.3 is the fractional edge cover number of the hypergraph $\{\{a,b\},\{b,c\},\{c,a\},\{c\}\}$; this number is 3/2. Any endomorphism of this open structure $(\mathbf{A},(a,b,c))$ has to fix $a$, $b$, and $c$, and so the only endomorphism is the identity map on $\{a,b,c\}$. This open structure is thus its own core. By Theorem 3.4, relative to $p_{1}$, a plan $p^{\prime}$ having the best possible intermediate degree has intermediate degree equal to $d:=C^{\mathbf{A}}_{\Sigma}\textup{-width}(\mathbf{A},(a,b,c))$; we compute this value as follows. The hypergraph $\mathbf{H}(\mathbf{A},(a,b,c))$ contains $\{a,b,c\}$ as an edge, so any tree decomposition of this hypergraph must contain a bag equal to $\{a,b,c\}$, and so any such tree decomposition $F$ must have $C^{\mathbf{A}}_{\Sigma}\textup{-width}(F)\geq C^{\mathbf{A}}_{\Sigma}(\{a,b,c\})$; as stated above, this latter value is $3/2$. On the other hand, one achieves a tree decomposition of this hypergraph by taking a one-vertex tree with bag $\{a,b,c\}$, and this tree decomposition $G$ has $C^{\mathbf{A}}_{\Sigma}\textup{-width}(G)=C^{\mathbf{A}}_{\Sigma}(\{a,b,c\})=3/2$. Thus the open structure has $C^{\mathbf{A}}_{\Sigma}\textup{-width}(\mathbf{A},(a,b,c))=3/2$.
Example A.3.
(output degree 1, best possible intermediate degree 3/2) Suppose we have a plan $p_{2}$ where $(\mathbf{A},(a,b))$ is a $p_{2}$-representation. An example of such a plan is $\pi_{1,2}\Join_{\theta}(E,E,E,E)$ where $\theta$ is as in Example A.2. By Proposition A.1, the output degree of $p_{2}$ is $C^{A}_{\Sigma}({a,b})$; by Proposition 2.3, this is the fractional edge cover number of a hypergraph with vertex set $\{a,b\}$ and that contains $\{a,b\}$ as an edge, so this output degree is 1. Any endomorphism of this open structure $(\mathbf{A},(a,b))$ has to fix $a$ and $b$, and no endomorphism can map $c$ to $a$ or $b$, since $(c,c)$ is an element of $E^{A}$ but $(a,a)$ and $(b,b)$ are not elements of $E^{A}$. Hence, the only endomorphism of this open structure is the identity map. This open structure is thus its own core. By Theorem 3.4, a plan $p^{\prime}$ having the best possible intermediate degree has intermediate degree equal to $C^{\mathbf{A}}_{\Sigma}\textup{-width}(\mathbf{A},(a,b))$; we compute this value as follows. The hypergraph $\mathbf{H}(\mathbf{A},(a,b))$ contains the edges $\{a,b\},\{a,c\},\{b,c\}$, and it is known that for any tree decomposition of a hypergraph and any clique appearing in a hypergraph, the tree decomposition must have a bag containing the clique. Hence any tree decomposition has a bag containing $\{a,b,c\}$, and by the analysis of Example A.2, we have $C^{\mathbf{A}}_{\Sigma}\textup{-width}(\mathbf{A},(a,b))=3/2$.
Example A.4.
(output degree 0, best possible intermediate degree 1) Suppose we have a plan $p_{3}$ where $(\mathbf{A},\epsilon)$ is a $p_{3}$-representation. An example of such a plan is $\pi_{\epsilon}\Join_{\theta}(E,E,E,E)$ where $\theta$ is as in Example A.2. By Proposition A.1, the output degree of $p_{3}$ is $C^{\mathbf{A}}_{\Sigma}(\epsilon)$, which is 0. Indeed, 0 is the output degree of any Boolean plan. The core of the open structure $(\mathbf{A},\epsilon)$ is readily verified to be $(\mathbf{B},\epsilon)$ where $\mathbf{B}$ is the structure with universe $\{c\}$ and the relation $E^{B}=\{(c,c)\}$; this is via the endomorphism of $(\mathbf{A},\epsilon)$ that maps each of $a$, $b$, and $c$ to $c$. By Theorem 3.4, a plan $p^{\prime}$ having the best possible intermediate degree has intermediate degree equal to $C^{\mathbf{B}}_{\Sigma}\textup{-width}(\mathbf{B},\epsilon)$. Since the hypergraph $\mathbf{H}(\mathbf{B},\epsilon)$ simply has vertex set $\{c\}$ and edge set $\{\{c\}\}$, this intermediate degree is $1$.
A.3 Proof of Proposition A.1
We show how the $p$-representation of a plan $p$ can be used to determine the output degree of the plan. To this end, we present two lemmas providing upper and lower bounds on the size of the result of evaluating an open structure on a structure. This lemma, providing an upper bound, is a reformulation of Theorem 4.4 of [10].
Lemma A.5.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $\mathbf{D}$ be a structure over $\sigma$ that satisfies $\Sigma$. Let $(\mathbf{A},\overline{a})$ be an open structure over $\sigma$. Let $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$. It holds that $|\mathsf{homs}(\mathbf{A},\{\overline{a}\},\mathbf{D})|\leq M_{\mathbf{D}}^{C_{\Sigma}^{\mathbf{B}}(\{\overline{b}\})}.$
We next present a complementary lemma that provides the matching lower bound and which again follows directly from Theorem 4.4 of [10].
Lemma A.6.
Let $\Sigma$ be a set of unary keys over signature $\sigma$. Let $\mathbf{A}$ be a structure over $\sigma$. For any open structure $(\mathbf{A},\overline{a})$, there exists a constant $c>0$ such that, with $\textrm{Chase}_{\Sigma}(\mathbf{A},\overline{a})=(\mathbf{B},\overline{b})$, there exists a sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures over $\sigma$ that satisfy $\Sigma$, with $M_{\mathbf{D}_{n}}$ unbounded, and where, for each $n\penalty 10000\ \geq\penalty 10000\ 1$, $|\mathsf{homs}(\mathbf{A},\{\overline{a}\},\mathbf{D}_{n})|\geq cM_{\mathbf{D}_{n}}^{C_{\Sigma}^{\mathbf{B}}(\{\overline{b}\})}.$
We now give the proof of the proposition. The existence of a $p$-representation follows from Theorem 4.6. Fix some structure $\mathbf{D}$ over $\sigma$. As $(\mathbf{A},\overline{a})$ is a $p$-representation, we have that $\mathsf{out}(p,\mathbf{D})=\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$. It follows from Lemma A.5 that $|\mathsf{out}(p,\mathbf{D})|\leq M_{\mathbf{D}}^{d}$, and so $p$ has output degree $\leq d$. By Lemma A.6, there is a constant $c>0$ and an infinite sequence of structures $(\mathbf{D}_{n})_{n\geq 1}$, with $M_{\mathbf{D}_{n}}$ unbounded, where $|\mathsf{out}(p,\mathbf{D}_{n})|\geq cM_{\mathbf{D}_{n}}^{d}$. We obtain that the plan $p$ has output degree $\geq d$, and can thus conclude by Lemma 4.1 that $p$ has output degree $d$.
Appendix B Proof of Section 2
B.1 Proof of Proposition 2.1
For the forward direction, let $h$ be a homomorphism as described; then, for any homomorphism $g$ from $\mathbf{A}^{\prime}$ to $\mathbf{D}$, the composition $g\circ h$ is a homomorphism from $\mathbf{A}$ to $\mathbf{D}$ where $g(\overline{a^{\prime}})=g(h(\overline{a}))$. For the backward direction, set $\mathbf{D}=\mathbf{A}^{\prime}$; then, we have that $\overline{a^{\prime}}\in\mathsf{homs}(\mathbf{A}^{\prime},\overline{a^{\prime}},\mathbf{D})$, implying that $\overline{a^{\prime}}\in\mathsf{homs}(\mathbf{A},\overline{a},\mathbf{D})$, and thus there exists a homomorphism from $\mathbf{A}$ to $\mathbf{A}^{\prime}$ sending $\overline{a}$ to $\overline{a^{\prime}}$.
Appendix C Proofs of Section 4
C.1 Proof of Lemma 4.1
Suppose that there existed a value $\epsilon>0$ such that a plan $p$ had output degree $\geq d$ and output degree $\leq(d-\epsilon)$. Then, there would be functions $g\in\Omega(M^{d})$ and $h\in O(M^{d-\epsilon})$ where $g(M_{\mathbf{D}_{n}})\leq|\mathsf{out}(p,\mathbf{D}_{n})|\leq h(M_{\mathbf{D}_{n}})$ for all structures $\mathbf{D}_{n}$ in a sequence $(\mathbf{D}_{n})_{n\geq 1}$ of structures with $M_{\mathbf{D}_{n}}$ unbounded, implying that $g\leq h$ on infinitely many values, a contradiction.
C.2 Proof of Lemma 4.2
The forward direction is immediate from the definitions. For the backward direction, we have that, for each subplan $q$ of $p$, there exists a function $h_{q}(M)\in O(M^{d})$ where, for all structures $\mathbf{D}$ over the signature of $p$, $|\mathsf{out}(q,\mathbf{D})|\leq h_{q}(M_{\mathbf{D}})$. The plan $p$ then has intermediate degree $\leq d$ via the function $\sum_{q}h_{q}$, where the sum is over all subplans $q$ of $p$.
C.3 Proof of Theorem 4.6
Before proceeding with the proof, we define an extended notion of representation that will facilitate our proving the existence of $p$-representations. Suppose that $\mathbf{A}$ is a structure, and $\approx$ is a binary relation on $\mathsf{dom}(\mathbf{A})$. We use the notation $\langle\approx\rangle$ to denote division by the equivalence relation $\approx^{*}$. So, for each element $a\in\mathsf{dom}(\mathbf{A})$, we use $a\langle\approx\rangle$ to denote the equivalence class of $\approx^{*}$ containing $a$. For each tuple $\overline{a}=(a_{1},\ldots,a_{k})$, we use $\overline{a}\langle\approx\rangle$ to denote $(a_{1}\langle\approx\rangle,\ldots,a_{k}\langle\approx\rangle)$. We use $\mathbf{A}\langle\approx\rangle$ to denote the structure with universe $\mathsf{dom}(\mathbf{A})\langle\approx\rangle$, defined as $\bigcup_{a\in\mathsf{dom}(\mathbf{A})}a\langle\approx\rangle$, and where for each symbol $R$, it holds that $R^{\mathbf{A}\langle\approx\rangle}=\{\overline{a}\langle\approx\rangle\penalty 10000\ |\penalty 10000\ \overline{a}\in R^{\mathbf{A}}\}$. We define an extended $p$-representation of a SPJ plan $p$ as a triple $(\mathbf{A},\overline{a},\approx)$ where $(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle)$ is a $p$-representation. A $p$-decomposition of an extended $p$-representation $(\mathbf{A},\overline{a},\approx)$ is defined as a $p$-decomposition of $(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle)$. We are now ready to start the actual proof of the theorem. We show, by induction on the structure of $p$, how to derive an extended $p$-representation $(\mathbf{A},\overline{a},\approx)$ and $p$-decomposition thereof which satisfies the containment property. We then have that $(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle)$ is a $p$-representation along with a $p$-decomposition. We prove the statement inductively over the structure of $p$. Suppose that $p=R$. Let $m$ be the arity of $R$. Define $\mathbf{A}$ as the structure with $A=\{a_{1},\ldots,a_{m}\}$, $R^{\mathbf{A}}=\{(a_{1},\ldots,a_{m})\}$, and $S^{\mathbf{A}}=\emptyset$ for each $S\in\sigma\setminus\{R\}$. Assign $\approx$ to be the empty set; we have that $\approx^{*}$ is the equality relation on $\mathsf{dom}(\mathbf{A})$. Letting $\overline{a}$ denote $(a_{1},\ldots,a_{m})$, we have that $(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle)$ is isomorphic to $(\mathbf{A},\overline{a})$. It is evident that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation. We give a $p$-decomposition of this representation as follows. Let $T$ be the tree with one vertex $t$, and define $\chi(t)=\{\overline{a}\}$. We then have that $(T,\chi)$ is a $p$-decomposition, via the mapping $\alpha$ with $\alpha(p)=t$ and the mapping $\beta$ with $\beta(p)=\overline{a}$. The containment property is, in this case, a direct consequence of the established fact that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation. Suppose that $p$ has the form $\pi_{j_{1},\ldots,j_{n}}(p_{0})$, where $p_{0}$ has arity $m$ and $j_{1},\ldots,j_{n}\in[m]$. By induction there exist an extended $p_{0}$-representation $(\mathbf{A},(a_{1},\ldots,a_{m}),\approx)$ and a $p_{0}$-decomposition $(T_{0},\chi_{0})$ thereof with root $r_{0}$ and mappings $\alpha_{0}$, $\beta_{0}$. Define $\overline{a}$ to be the tuple $(a_{j_{1}},\ldots,a_{j_{n}})$; we then have that for any structure $\mathbf{D}$ over $\sigma$: $\mathsf{homs}(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle,\mathbf{D})=\pi_{j_{1},\ldots,j_{n}}(\mathsf{homs}(\mathbf{A}\langle\approx\rangle,(a_{1},\ldots,a_{m})\langle\approx\rangle,\mathbf{D}))$ and thus that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation. We obtain a $p$-decomposition $(T,\chi)$ of this representation by extending $(T_{0},\chi_{0})$ as follows: create a root node $r$ whose unique child is $r_{0}$, and let $\chi(t)=\chi_{0}(t)$ for every $t\in T_{0}$ and define $\chi(r)=\{\overline{a}\langle\approx\rangle\}$. Extend the mappings $\alpha_{0}$, $\beta_{0}$ to mappings $\alpha,\beta$ by defining $\alpha(p)=r$ and $\beta(p)=\overline{a}\langle\approx\rangle$. The containment property on each subplan of $p_{0}$ holds as a direct consequence of it having held on the $p_{0}$-decomposition. When we consider $p$ (as a subplan of $p$), the containment property holds due to $(\mathbf{A},\overline{a},\approx)$ being an extended $p$-representation. Suppose that $p$ has the form $\Join_{\theta}(p_{1},\ldots,p_{\ell})$, where each $p_{i}$ has arity $m_{i}$, for $i\in[\ell]$, and $p$ has arity $m=m_{1}+\cdots+m_{\ell}$. By induction, there exists (for each $i$) an extended $p_{i}$-representation $(\mathbf{A}_{i},\overline{t_{i}},\approx_{i})$. We may assume without loss of generality that the universes $\mathsf{dom}(\mathbf{A}_{i})$ are pairwise disjoint (if they are not, their elements can be renamed to achieve this). We let $\overline{a}=(a_{1},\ldots,a_{m})$ denote the tuple $\overline{t_{1}}\ldots\overline{t_{\ell}}$, that is, the tuple obtained by concatenating the entries of the tuples $\overline{t_{i}}$, $i\in[\ell]$, in order. Define $\approx_{0}$ as the relation $\bigcup_{i=1}^{\ell}\approx_{i}$, i.e., collecting all previously known equivalences, and define $\approx$ as the relation $\approx_{0}\cup\bigcup_{(j=k)\in\theta}\{(a_{j},a_{k})\}$, i.e., adding the equivalences newly introduced by $\theta$. We define $\mathbf{A}$ as $\mathbf{A}_{1}\cup\cdots\cup\mathbf{A}_{\ell}$. Then we have that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation, since for any structure $\mathbf{D}$ over $\sigma$: $\mathsf{homs}(\mathbf{A}\langle\approx\rangle,\overline{a}\langle\approx\rangle,\mathbf{D})$ $=\sigma_{\theta}(\mathsf{homs}(\mathbf{A}\langle\approx_{0}\rangle,\overline{a}\langle\approx_{0}\rangle,\mathbf{D}))$ $=\sigma_{\theta}(\mathsf{homs}(\mathbf{A}_{1}\langle\approx_{1}\rangle,\overline{t_{1}}\langle\approx_{1}\rangle,\mathbf{D})\times\cdots\times\mathsf{homs}(\mathbf{A}_{\ell}\langle\approx_{\ell}\rangle,\overline{t_{\ell}}\langle\approx_{\ell}\rangle,\mathbf{D}))$ $=\sigma_{\theta}(\mathsf{out}(p_{1},\mathbf{D})\times\cdots\times\mathsf{out}(p_{\ell},\mathbf{D}))$ $=\mathsf{out}(p,\mathbf{D}).$ Here, the first equality comes from our definition of $\approx$ from $\approx_{0}$, and the fact that each pair in $\approx_{0}$ has elements contained entirely in a universe $\mathsf{dom}(\mathbf{A}_{i})$; the second equality follows from this fact and our definitions of $\overline{a}$ and $\approx_{0}$; the third equality follows from our assumption that each $(\mathbf{A}_{i},\overline{t_{i}},\approx_{i})$ is an extended $p_{i}$-representation; the last equality follows from the definition of $\mathsf{out}$. By induction, each given extended $p_{i}$-representation has a $p_{i}$-decomposition $E_{i}=(T_{i},\chi_{i})$, where $r_{i}$ is the root node of $T_{i}$, with associated mappings $\alpha_{i},\beta_{i}$. We assume without loss of generality that the vertex sets $V(T_{i})$ are pairwise disjoint (if not, they can be renamed to achieve this). We provide a $p$-decomposition as follows. Combining the trees $T_{i}$, create a rooted tree $T$ with one extra vertex $r$ whose children are $r_{1},\ldots,r_{\ell}$. Define $\alpha:\mathsf{subplans}(p)\to V(T)$ to be the mapping that extends each mapping $\alpha_{i}$ and where $\alpha(p)=r$; define $\beta:\mathsf{subplans}(p)\to\mathsf{dom}(\mathbf{A})\langle\approx\rangle^{*}$ to be the mapping with $\beta(p)=\overline{a}\langle\approx\rangle$, and where, for each subplan $q$ of a plan $p_{i}$, when $\beta_{i}(q)=(c_{1},\ldots,c_{m_{i}})$, $\beta(q)$ is defined as the tuple obtained by further dividing each entry by $\approx^{*}$, that is, as the tuple $(c^{\prime}_{1},\ldots,c^{\prime}_{m_{i}})$ where each $c^{\prime}_{j}$ is the unique $\approx^{*}$-equivalence class containing $c_{j}$; such a class exists, since each $c_{j}$ is a $\approx_{i}^{*}$-equivalence class, and $\approx_{i}^{*}$ is contained in $\approx^{*}$ (due to $\approx_{i}$ being contained in $\approx$). Define $\chi(r)=\{\overline{a}\langle\approx\rangle\}$. For each subplan $q$ of $p$ with $q\neq p$, define $\chi({\alpha(q)})=\{\beta(q)\}$. It is straightforward to verify that $(T,\chi)$ is a rooted tree decomposition as required by the definition of $p$-decomposition with mappings $\alpha$ and $\beta$ as defined above. It remains to show that the $p$-decomposition $(T,\chi)$ satisfies the containment property. We showed that $(\mathbf{A},\overline{a},\approx)$ is an extended $p$-representation, so it holds in the case that $q=p$. Suppose that $q$ is a subplan of $p$ with $q\neq p$. Then $q$ is a subplan of a plan $p_{j}$, for some $j\in[\ell]$. Since $(\mathbf{A}_{j},\overline{t_{j}},\approx_{j})$ is an extended $p_{j}$-representation, we have for any structure $\mathbf{D}$ over $\sigma$ that $\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},\beta_{j}(q),\mathbf{D}).$ Let us make some observations. First, we have $\alpha(q)=\alpha_{i}(q)$, and so $\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$ is a substructure of $\mathbf{A}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$. We can write $\beta_{j}(q)$ as a tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle$ where $a^{\prime}_{1},\ldots,a^{\prime}_{k}\in\mathsf{dom}(\mathbf{A}_{j})$. By our definition of $\beta(q)$ from $\beta_{j}(q)$, we have $\beta(q)=(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$. We have $\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},\beta_{j}(q),\mathbf{D})$ $=\mathsf{homs}(\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}},(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle,\mathbf{D})$ $\supseteq\mathsf{homs}(\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E},(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle,\mathbf{D}).$ The equality follows from the above observations on $\beta_{j}(q)$. The containment can be explained as follows. We have that $\approx_{j}$ is a subset of $\approx$, and, on nodes below $\alpha(q)$, the tree decomposition $E$ is obtained from the tree decomposition $E_{j}$ by further dividing each element in a set associated via $\chi_{j}$ by $\approx^{*}$. We thus have a canonical homomorphism from $\mathbf{A}_{j}\langle\approx_{j}\rangle_{\leq\alpha_{j}(q),E_{j}}$ to $\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E}$, namely, the mapping that further divides each element of the first structure by $\approx^{*}$. Since this homomorphism maps the tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx_{j}\rangle$ to the tuple $(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$, the containment holds by Proposition 2.1. By chaining together the established relationships and using $\beta(q)=(a^{\prime}_{1},\ldots,a^{\prime}_{k})\langle\approx\rangle$, we conclude $\mathsf{out}(q,\mathbf{D})\supseteq\mathsf{homs}(\mathbf{A}\langle\approx\rangle_{\leq\alpha(q),E},\beta(q),\mathbf{D})$, as required.
C.4 Proof of Theorem 4.7
The heart of the proof can be derived from argumentation in Section 4 of [10]. As we draw on this section, we employ the terminology of [10], and refer to their argumentation. In particular, their result refers to conjunctive queries and functional dependencies. The former are known to be equivalent to SPJ queries (and are in syntax close to our $p$-representations), while the latter generalize key constraints. In the following arguments, as in [10], for elements $X_{i},X_{j}\in\mathsf{dom}(\mathbf{C})$, we use the notation $X_{i}\rightarrow X_{j}$ to say that element $X_{j}$ is functionally determined by $X_{i}$ via some functional dependency. In Section 4 of [10], it is explained how to transform a conjunctive query $Q_{c}$, to which the chase has been applied, to a conjunctive query $Q^{\prime}$ over a modified set of relations that has no functional dependencies. Let us for the moment put aside the output relation and output variables, and consider the action of this transformation on the input relations and variables. The transformation considers functional dependencies of the form $X_{i}\rightarrow X_{j}$ over pairs of input variables, and iteratively removes them; a removal is performed by adding entailed functional dependencies, and by expanding each atom including $X_{i}$ to a new atom where $X_{j}$ is added (if not already present). For each functional dependency $X_{i}\rightarrow X_{j}$, we can maintain a binary relation $S_{ij}$ whose first coordinate is a key for the relation, and such that adding $S_{ij}(X_{i},X_{j})$ to the query atoms would not change the query answers; this holds initially by taking $S_{ij}$ to be an appropriate projection of the relation of an atom mentioning both $X_{i}$ and $X_{j}$, and then when an entailed dependency $X_{k}\rightarrow X_{j}$ is added, due to dependencies $X_{k}\rightarrow X_{i}$ and $X_{i}\rightarrow X_{j}$, we can derive $S_{kj}$ (if needed) by taking the natural join of $S_{ki}$ and $S_{ij}$ and then projecting the middle coordinate. It is readily verified that the sizes of these relations $S_{ij}$ never exceeds the size of any input relation. When an atom including a variable $X_{i}$ is expanded to include the variable $X_{j}$, a new relation for the atom can be computed by joining the original relation with the relation $S_{ij}$. Since each relation $S_{ij}$ expresses a functional dependency, this never increases the size of any relation. The result of the transformation, applied to the input atoms of a conjunctive query, is thus a new set of input atoms where every new relation can be derived from the original relations via a well-behaved SPJ plan of intermediate degree $1$, where every new relation has size bounded above by an original relation, and where the answers (over all variables) to the new set of atoms is the same as the answers to the original set. Consider the application of the transformation to the input atoms of the conjunctive query corresponding to $\mathbf{C}$. Let $\mathbf{C}^{\prime}$ be the relational structure representing the new set of atoms. It is moreover shown in Lemma 4.7 of [10] that the color number of an overall query (including input variables) is not changed by this transformation, and this implies that $C^{\mathbf{C}^{\prime}}_{\Sigma}(S)\leq C^{\mathbf{C}^{\prime}}_{\Sigma}(S^{\prime})=C^{\mathbf{C}}_{\Sigma}(S)$ for any set $S$ of variables and its expansion $S^{\prime}$ after applying the transformation. Let us then consider $(\mathbf{C}^{\prime},\overline{c})$; due to the just-given inequality, we have $C_{\emptyset}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})\leq C_{\Sigma}^{\mathbf{C}}\textup{-width}(\mathbf{C},\overline{c})$, and it thus suffices to establish the result for the empty set of unary keys $\Sigma^{\prime}=\emptyset$ along with $(\mathbf{C}^{\prime},\overline{c})$, in place of $\Sigma$ with $(\mathbf{C},\overline{c})$. To establish the result for $(\mathbf{C}^{\prime},\overline{c})$ in the absence of keys, it suffices to compute a tree decomposition $F$ of $(\mathbf{C}^{\prime},\overline{c})$ where $C_{\Sigma^{\prime}}^{\mathbf{C}^{\prime}}\textup{-width}(F)$ is equal to $C_{\Sigma^{\prime}}^{\mathbf{C}^{\prime}}\textup{-width}(\mathbf{C}^{\prime},\overline{c})$, which can be done in exponential time, and then compute a plan that materializes the answers to each bag, where the answers for a bag with variables $S$ would be considered to be the join of the set of atoms obtained from $\mathbf{C}^{\prime}$’s conjunctive query by taking each atom, and projecting onto $S$. These different sets of answers can then be combined via joins and projects in a standard way, following the structure of the tree decomposition. But to find a plan that does this materialization for a bag with variables $\{a_{1},\ldots,a_{n}\}$, we simply follow the proof of Theorem 6 of [2], which explains how to construct a join-project plan for any pure join plan; in doing so, they define plans $\phi_{1},\ldots,\phi_{n}$; each of these plans is implementable as a well-behaved one, since they iteratively introduce one variable at a time: in their notation, the plan $\phi_{i}$, which can be viewed as a multiway join, concerns variables $\{a_{1},\ldots,a_{i}\}$.
C.5 Proof of Lemma 4.9
By the definition of $C_{\Sigma}^{\mathbf{C}}\textup{-width}$ of the open structure $(\mathbf{C},\overline{c})$, there exists a vertex $t_{0}$ of the given tree decomposition such that $C_{\Sigma}^{\mathbf{C}}\textup{-width}(\chi(t_{0}))\geq d$. Let $\overline{\chi(t_{0})}$ be the tuple of elements in $\chi(t_{0})$. Invoke Lemma A.6 on the structure $\mathbf{C}$ and the open structure $(\mathbf{C},\overline{\chi(t_{0})})$ to obtain a constant $c^{\prime}>0$ and structures $(\mathbf{D}_{n})_{n\geq 1}$ such that $|\mathsf{homs}(\mathbf{C},\chi(t_{0}),\mathbf{D}_{n})|\geq c^{\prime}M_{\mathbf{D}_{n}}^{C_{\Sigma}^{\mathbf{C}}(\chi(t_{0}))}$.
C.6 Proof of Lemma 4.10
We show how to pass from any map in $\mathsf{homs}(\mathbf{C},S\penalty 10000\ \cap\penalty 10000\ \mathsf{dom}(\mathbf{C}),\mathbf{D})$ to a map in $\mathsf{homs}(\mathbf{B},S,\mathbf{D})$, and then explain why the passage is injective, which suffices. Suppose that $g\in\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\mathbf{D})$ and $h$ is a retraction from $\mathbf{B}$ to $\mathbf{C}$. Then there exists an extension $g^{\prime}$ of $g$ that is in $\mathsf{homs}(\mathbf{C},\mathsf{dom}(\mathbf{C}),\mathbf{D})$, from which it follows that the composition $g^{\prime}\circ h$ is in $\mathsf{homs}(\mathbf{B},\mathsf{dom}(\mathbf{B}),\mathbf{D})$ (in writing this composition, $h$ is applied first). It follows that the restriction $(g^{\prime}\circ h)\upharpoonright S$ is in $\mathsf{homs}(\mathbf{B},S,\mathbf{D})$. This passage is injective since if two mappings in $\mathsf{homs}(\mathbf{C},S\cap\mathsf{dom}(\mathbf{C}),\mathbf{D})$ differ on an element $c$ of $S\cap\mathsf{dom}(\mathbf{C})$, the two mappings that they are passed to will also differ on $c$, due to the assumption that $h$ acts as the identity on $C$.
Appendix D Proofs of Section 5
D.1 Proof of Theorem 5.1
We first show how to pass from an SPJU plan $p$ to a sequence of SPJ plans $(p_{i})$ and a sequence of maps $(\alpha_{i})$ satisfying the first two conditions. We show this by induction on the structure of $p$. When $p=R$, we take $p_{1}=R$ and define $\alpha_{1}$ by $\alpha_{1}(p_{1})=p$. When $p=\pi_{j_{1},\ldots,j_{n}}(p^{\prime})$, by induction, for the SPJU plan $p^{\prime}$ we have SPJ plans $p^{\prime}_{1},\ldots,p^{\prime}_{m^{\prime}}$ with maps $\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{m^{\prime}}$. For each $i=1,\ldots,m^{\prime}$, define $p_{i}=\pi_{j_{1},\ldots,j_{n}}(p^{\prime}_{i})$. It is straightforwardly verified that $p$ is semantically equivalent to $p_{1}\cup\cdots\cup p_{m}$. For each $i=1,\ldots,m^{\prime}$, define $\alpha_{i}:\mathsf{subplans}(p_{i})\to\mathsf{subplans}(p)$ to be the unique extension of $\alpha^{\prime}_{i}$ such that $\alpha_{i}(p_{i})=p$; this satisfies the desired property due to the aforementioned semantic equivalence, and induction. When $p=\Join_{\theta}(p_{1},\ldots,p_{\ell})$, by induction, we have for each $p_{i}$ a sequence of plans $p_{i,1},\ldots,p_{i,m_{i}}$ with maps $\alpha_{i,1},\ldots,\alpha_{i,m_{i}}$. For each index sequence $j_{1},\ldots,j_{\ell}$ with $1\leq j_{i}\leq m_{i}$ (for each $i\in[\ell]$), define $q_{j_{1},\ldots,j_{\ell}}$ as the plan $\Join_{\theta}(p_{1,j_{1}},\ldots,p_{\ell,j_{\ell}})$. It is straightforward to verify that $p$ is semantically equivalent to the union $(\cup)$ over all of these plans $q_{j_{1},\ldots,j_{\ell}}$. To define the map $\alpha_{j_{1},\ldots,j_{\ell}}$ corresponding to the plan $q_{j_{1},\ldots,j_{\ell}}$, we combine the maps $\alpha_{1,j_{1}},\ldots,\alpha_{\ell,j_{\ell}}$, and extend the combination to map $q_{j_{1},\ldots,j_{\ell}}$ to $p$; this satisfies the desired property due to the aforementioned semantic equivalence, and induction. When $p=p^{\prime}\cup p^{\prime\prime}$, by induction, for the SPJU plan $p^{\prime}$ we have SPJ plans $p^{\prime}_{1},\ldots,p^{\prime}_{m^{\prime}}$, and for the SPJU plan $p^{\prime\prime}$ we have SPJ plans $p^{\prime\prime}_{1},\ldots,p^{\prime\prime}_{m^{\prime\prime}}$. We simply take the concatenation of these two SPJ plan sequences, and retain the associated maps. It is straightforwardly verified that the first two conditions hold. To prove the theorem, consider the following algorithm. Given an SPJU plan $p$, it first uses the passage just given to compute a sequence of SPJ plans $p_{1},\ldots,p_{m}$ with associated maps. It is straightforward to verify that this can be done in singly exponential time: in the case $p=\Join_{\theta}(p_{1},\ldots,p_{\ell})$, the number of plans for $p$ is the product (over $i=1,\ldots,\ell$) of the number of plans for $p_{i}$, and in the case $p=p^{\prime}\cup p^{\prime\prime}$, the number of plans for $p$ is the sum of the number of plans for $p^{\prime}$ and $p^{\prime\prime}$; it follows by a straightforward induction that the total number of plans for any subplan $p^{+}$ is at most $2$ raised to the number of leaves below $p^{+}$, and is thus singly exponential. From Theorem 4.6, for each plan $p_{i}$, the algorithm computes a $p_{i}$-representation $(\mathbf{A}_{i},\overline{a_{i}})$. Next, it continually checks to see if there is a pair $(i,j)$ of indices where there exists a homomorphism from $(\mathbf{A}_{i},\overline{a_{i}})$ to $(\mathbf{A}_{j},\overline{a_{j}})$; checking for such a homomorphism requires singly exponential time. If there is such a pair, the algorithm arbitrarily selects one such pair $(i,j)$, and removes the plan $p_{j}$ (and its associated map) from the sequence; this removal preserves property (1), due to Proposition 2.1. When this process terminates, all three properties of the theorem are satisfied.
D.2 Proof of Theorem 3.3
Given an SPJU plan $p$, the algorithm first uses the algorithm of Theorem 5.1 to obtain, in exponential time, a sequence of SPJ plans $p_{1},\ldots,p_{m}$ with open structures $(\mathbf{A}_{1},\overline{a_{1}}),\ldots,(\mathbf{A}_{m},\overline{a_{m}})$ such that $p$ is semantically equivalent (and thus $\Sigma$-semantically equivalent) to $p_{1}\cup\cdots\cup p_{m}$. For each index $i$, we have that $(\mathbf{A}_{i},\overline{a_{i}})$ is a $p_{i}$-representation. The algorithm then does the following for each plan $p_{i}$: via Theorem 3.4 we obtain, in exponential time, a value $d_{i}$ and a plan $p^{\prime}_{i}$ such that $p^{\prime}_{i}$ is $\Sigma$-semantically equivalent to $p_{i}$ and $p^{\prime}_{i}$ has intermediate degree $d_{i}$, which is the best possible intermediate degree. Finally, the algorithm outputs the value $d=\max_{i\in[m]}d_{i}$ and the plan $p^{\prime}=p^{\prime}_{1}\cup\cdots\cup p^{\prime}_{m}$. Since each plan $p^{\prime}_{i}$ has intermediate degree $\leq d$, we have that $p^{\prime}$ has intermediate degree $\leq d$. Moreover, choosing $k$ to be an index where $d_{k}=d$, we have that $p^{\prime}_{k}$ does not have (for any $\epsilon>0$) intermediate degree $\leq(d-\epsilon)$, implying that $p^{\prime}$ has intermediate degree $d$. To complete the proof, it suffices to argue that, for any SPJU plan $q^{\prime}$ $\Sigma$-semantically equivalent to $p^{\prime}$, and any $\epsilon>0$, it holds that $q^{\prime}$ does not have intermediate degree $\leq(d-\epsilon)$. Let $k$ be an index where $d=d_{k}$. By Theorem 5.1 applied to $q^{\prime}$, there exists a sequence $q^{\prime}_{1},\ldots,q^{\prime}_{n}$ of SPJ plans, along with representations $(\mathbf{B}_{1},\overline{b_{1}}),\ldots,(\mathbf{B}_{n},\overline{b_{n}})$ satisfying the conditions of the theorem. We claim that there exists an index $\ell\in[n]$ where $q^{\prime}_{\ell}$ is $\Sigma$-semantically equivalent to $p^{\prime}_{k}$. Let $(\mathbf{C}_{k},\overline{c_{k}})$ be a core of the chase of $(\mathbf{A}_{k},\overline{a_{k}})$. Since $C_{k}$ satisfies $\Sigma$, we have $\overline{c_{k}}\in\mathsf{out}(p^{\prime}_{k},\mathbf{C}_{k})\subseteq\mathsf{out}(p^{\prime},\mathbf{C}_{k})=\mathsf{out}(q^{\prime},\mathbf{C}_{k})$, implying that there exists an index $\ell$ with $\overline{c_{k}}\in\mathsf{out}(q^{\prime}_{\ell},\mathbf{C}_{k})$. So, there exists a homomorphism from $(\mathbf{B}_{\ell},\overline{b_{\ell}})$ to $(\mathbf{C}_{k},\overline{c_{k}})$. By a symmetric argument, there exists an index $k^{\prime}$ such that there is a homomorphism from $(\mathbf{C}_{k^{\prime}},\overline{c_{k^{\prime}}})$ to $(\mathbf{B}_{\ell},\overline{b_{\ell}})$. By condition (3) of Theorem 5.1, since we have a homomorphism from $(\mathbf{C}_{k^{\prime}},\overline{c_{k^{\prime}}})$ to $(\mathbf{C}_{k},\overline{c_{k}})$, we have $k^{\prime}=k$, and so $q^{\prime}_{\ell}$ is $\Sigma$-semantically equivalent to $p^{\prime}_{k}$. By Theorem 4.8 and Lemma 4.1, $q^{\prime}_{\ell}$ does not have (for any $\epsilon>0$) output degree $\leq(d-\epsilon)$, and thus it does not have intermediate degree $\leq(d-\epsilon)$ implying by condition (2) of Theorem 5.1 that $q^{\prime}$ does not have (for any $\epsilon>0$) intermediate degree $\leq(d-\epsilon)$.